### v1

# Transverse Momentum-Dependent Heavy-Quark Fragmentation at Next-to-Leading Order

Abstract
The transverse momentum-dependent fragmentation functions (TMD FFs) of heavy (bottom and charm) quarks, which we recently introduced, are universal building blocks that enter predictions for a large number of observables involving final-state heavy quarks or hadrons. They enable the extension of fixed-order subtraction schemes to quasi-collinear limits, and are of particular interest in their own right as probes of the nonperturbative dynamics of hadronization. In this paper we calculate all TMD FFs involving heavy quarks and the associated TMD matrix element in heavy-quark effective theory (HQET) to next-to-leading order in the strong interaction. Our results confirm the renormalization properties, large-mass, and small-mass consistency relations predicted in our earlier work. We also derive and confirm a prediction for the large-$z$ behavior of the heavy-quark TMD FF by extending, for the first time, the formalism of joint resummation to capture quark mass effects in heavy-quark fragmentation. Our final results in position space agree with those of a recent calculation by another group that used a highly orthogonal organization of singularities in the intermediate momentum-space steps, providing a strong independent cross check. As an immediate application, we present the complete quark mass dependence of the energy-energy correlator (EEC) in the back-to-back limit at $\mathcal{O}(\alpha_{s})$.
1 Introduction
The fragmentation of heavy (bottom or charm) quarks into the experimentally observed heavy meson and baryon states is of particular interest because the mass of the quark imprints as a perturbative scale on the otherwise nonperturbative process of hadronization. In ref. [1], we initiated the study of transverse momentum-dependent (TMD) fragmentation functions (FFs) for the formation of heavy hadrons from a parent heavy quark. Our work generalized the well-studied case where only the collinear momentum fraction of the hadron is resolved [2, 3, 4, 5, 6] (see refs. [7] for a recent precision study on data), and is part of a larger, ongoing research effort to understand differential jet and fragmentation functions involving heavy quarks, heavy hadrons, and heavy-quark bound states [8, 9, 1, 10, 11, 12, 13, 14]. In this context, TMD FFs are unique in the wealth of information they can provide on the hadronization process, essentially offering a full three-dimensional view of the fragmentation cascade. The main goal of the present paper is to evaluate all unpolarized heavy-quark TMD FFs at perturbative transverse momenta to complete next-to-leading order (NLO) in the strong coupling, with the aim of improving the baseline perturbative precision for studying their rich nonperturbative structure.
Importantly, the factorization paradigm ensures that the heavy-quark TMD FFs appear as universal building blocks across predictions for a large number of processes involving final-state heavy quarks. In ref. [1], we explicitly considered their phenomenology in heavy-hadron pair production in the back-to-back limit in $e^{+}e^{-}$ collisions and in (polarized) semi-inclusive DIS at the future Electron-Ion Collider [15], cf. ref. [16] for recent dedicated projections for polarized collinear parton distribution functions (PDFs) and FFs involving heavy quarks at the EIC. However, the applicability of heavy-quark TMD FFs is much wider: Applications of immediate phenomenological interest range from $Z$$+$hadron [17] and dihadron [18] azimuthal decorrelations in $pp$ and $pA$ collisions to hadron-in-jet transverse-momentum distributions [19, 20] and energy-energy correlators in the back-to-back limit in $e^{+}e^{-}$ [21], $ep$ and $eA$ [22, 23], and $pp$ collisions [24, 25]. In all of these cases, factorization formulas involving TMD FFs have been derived, and the NLO results we present in this paper enable one to fully account for the effect of quark masses and accurately capture their highly nontrivial interplay with transverse momenta in all of these processes.
Figure -1079: The probability density $\mathrm{d}\mathcal{P}_{Q\bar{Q}\to(Qg)\bar{Q}}$ for a gluon to split off a heavy quark in the quasi-collinear limit, which coincides with the heavy-quark TMD FF $d_{1\,Q/Q}$ for finite collinear gluon momentum fraction $\bar{z}>0$ and transverse momentum $k_{T}>0$. On the rear walls of the plot we show the integrals of $d_{1\,Q/Q}$ over one of $k_{T}$ or $z$ as predicted by the NLO calculation we perform in this paper, showcasing that the TMD regularization and renormalization procedure renders them finite despite the intricate singular behavior of the integrand. The same is true for the total integral over both $z$ and $k_{T}$, which is not shown in the plot, but at the values of the scales chosen here is equal to $\frac{\alpha_{s}C_{F}}{4\pi}\bigl{(}5-\frac{\pi}{2}-\frac{2\pi^{2}}{3}\bigr{)}$. For numerical illustration, we take $\alpha_{s}(\mu=m)=0.2$.
The heavy-quark TMD FF is intimately connected with the universal gluon splitting probability off a heavy-quark in the quasi-collinear limit [26, 27], e.g. after heavy-quark pair production has taken place,
$\frac{\mathrm{d}\mathcal{P}_{Q\bar{Q}\to(Qg)\bar{Q}}}{\mathrm{d}z\,\mathrm{d}(k_{T}^{2})}=\frac{\alpha_{s}C_{F}}{2\pi}\,\frac{k_{T}^{2}z^{2}(1+z^{2})+m^{2}(1-z)^{4}}{(1-z)\bigl{[}k_{T}^{2}z^{2}+m^{2}(1-z)^{2}\bigr{]}^{2}}\,,$ (1.1)
where $m$ is the heavy-quark mass, $k_{T}$ is the transverse momentum of the gluon with respect to the heavy quark, and $z$ is the collinear momentum fraction retained by the heavy quark. The quasi-collinear splitting amplitude is shown in a three-dimensional plot as a function of $k_{T}$ and the momentum fraction $\bar{z}\equiv 1-z$ carried by the gluon in figure -1079. Famously, and in contrast to the behavior of light partons, the splitting amplitude is suppressed for energetic gluons with large momentum fractions $\bar{z}\gtrsim 0.2$, an effect known as the dead cone [28]. The splitting probability is strongly peaked at small $\bar{z}$ and small $k_{T}\lesssim m$. Unlike for light partons, it eventually turns over for any finite value of $\bar{z}>0$ and vanishes (near the front right edge of the plot) as $k_{T}\to 0$, yet the simultaneous limit $\bar{z},k_{T}\to 0$ diverges. A key challenge we address in this paper is how to consistently extend this complicated singular behavior of the quasi-collinear splitting probability into the limits $k_{T}\to 0,\bar{z}\to 0$ where the gluon is unresolved. The right framework for doing so is – indeed – TMD factorization [29], which provides precise field-theoretic prescriptions to isolate, regulate, and renormalize the associated divergences, and it is precisely in this sense that the heavy-quark TMD FF is the formal generalization of eq. (1.1) to $\bar{z}\to 0$ and/or $k_{T}\to 0$ in a way that is manifestly finite in four spacetime dimensions. As a demonstration, we include on the rear walls of the three-dimensional plot in figure -1079 the integral of the heavy-quark TMD FF over one of $z$ or $k_{T}\leq m$, respectively, as predicted by our calculations in this paper, showcasing that they are indeed individually finite. Similarly, the total integral over both variables $0\leq z\leq 1$ and $0\leq k_{T}\leq m$ (and thus any integral over a subset of the plane) is also finite in the usual distributional sense, and for the indicated values of the scales simply evaluates to $\frac{\alpha_{s}C_{F}}{4\pi}\bigl{(}5-\frac{\pi}{2}-\frac{2\pi^{2}}{3}\bigr{)}$. We stress that isolating this singular behavior at fixed order in terms of a universal and finite heavy-quark TMD FF is not an academic exercise, but instead is of immediate interest to extend fixed-order subtraction schemes like that of $q_{T}$ subtractions [30] and its multi-jet generalizations [31, 32, 33, 34] to quasi-collinear limits near final-state heavy quarks. Such extensions, which can potentially yield large gains in numerical efficiency, are likewise an active topic of current research, see e.g. [35, 36, 37, 38].
The remainder of this paper is structured as follows: Section 2 contains a review of the theoretical framework introduced in ref. [1] to describe heavy-quark fragmentation in the transverse plane. In section 3, we calculate all heavy-quark TMD FFs to next-to-leading order in the strong coupling, and present the results in both momentum and position space. In section 4, we perform an analogous calculation for the effective TMD matrix elements that arise in boosted heavy-quark effective theory (bHQET) in the limit where the transverse momentum is small compared to the mass. We employ these results and known one-loop results in the literature to verify the expected small and large-mass consistency relations in section 5. Section 6 is dedicated to an extension of the joint resummation formalism to massive quark fragmentation, which explains the surprising simplicity of our results at large $z$. As an immediate, first application of our results, we present the complete mass dependence of the energy-energy correlator in the back-to-back limit in section 7. In section 8, we compare our main results to those of ref. [13], where an independent calculation of the one-loop heavy-quark TMD FF was recently performed. We find full agreement at the level of the final position-space results, but stress the differences in how the intermediate momentum-space calculations are organized (amounting to a strong independent cross check) and – on the phenomenological side – the differences in how the $z$ dependence is treated in the bHQET limit. We conclude in section 9.
2 Theoretical framework
2.1 Heavy-quark TMD FFs
We study the fragmentation of a heavy quark $Q$ with pole mass $m\gg\Lambda_{\mathrm{QCD}}$ into an unpolarized heavy hadron $H$ and additional fragmentation products $X$. We work in QCD with $n_{f}=n_{\ell}+1$ flavors, where $n_{\ell}$ is the number of massless quark flavors. We work in the “hadron frame” for fragmentation [29] where $P_{H}^{\mu}=(P_{H}^{-},M_{H}^{2}/P_{H}^{-},0)$, with $M_{H}$ the hadron mass, and $P_{H,\perp}=0$ by definition. (We refer to Appendix A.1 for our lightcone conventions.)
In ref. [1] we showed that up to a factor of the total probability for $Q$ to fragment into $H$, the distribution differential in $k_{T}=P_{X,T}\sim M_{H}$ and the lightcone momentum fraction $z_{H}$ of the original quark retained by $H$ is governed by a new perturbative matching coefficient, the partonic heavy-quark TMD FF $d_{1\,Q/Q}(z,k_{T},\mu,\zeta)$,
$D_{1\,H/Q}(z_{H},b_{T},\mu,\zeta)=d_{1\,Q/Q}(z_{H},b_{T},\mu,\zeta)\,\chi_{H}+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{m}\Bigr{)}+\mathcal{O}(\Lambda_{\mathrm{QCD}}b_{T})\,.$ (2.1)
Making all regulators explicit, the formal definition of $d_{1\,Q/Q}$ at the bare level reads
$d_{1\,Q/Q}^{\mathrm{bare}}(z,b_{T},\epsilon,\eta,\zeta/\nu^{2})$ $=\frac{1}{2z^{1-2\epsilon}N_{c}}\int\!\frac{\mathrm{d}b^{+}}{4\pi}\,e^{\mathrm{i}b^{+}(p^{-}/z)/2}$ (2.2)
[…]
where we work in $d=4-2\epsilon$ spacetime dimensions, $N_{c}$ is the number of colors, $\operatorname{Tr}$ ($\operatorname{tr}$) indicates a trace over color (spin), and
X
subscript
[…]
indicates a sum over all possible partonic final states combined with an integral over their phase space. The fields in the first matrix element are evaluated at a spacetime position $b\equiv(0,b^{+},b_{\perp})$ with $b_{\perp}$ Fourier conjugate to $k_{\perp}$, and $b_{T}^{2}\equiv-b_{\perp}^{2}$ and $k_{T}^{2}=-k_{\perp}^{2}$, respectively. The heavy quark in the external state carries momentum $p^{\mu}=(p^{-},m^{2}/p^{-},0)$, i.e., the above definition is equal to the hadron-level definition of the heavy-quark TMD FF $D_{1\,H/Q}$, but with $H$ replaced by $Q$ itself in the external state and restricting to partonic final states $X$ [1]. The definition in eq. (2.2) is given in terms of the heavy-quark field renormalized on shell,
$\psi^{\mathrm{bare}}_{Q}(x)=Z_{\psi,\mathrm{OS}}^{1/2}(m,\mu,\epsilon)\,\psi_{Q}(x)\,,$ (2.12)
and in terms of Wilson lines $W_{\eta}(x)$ defined as anti-path ordered exponentials of gauge fields extending to positive infinity along the lightcone direction $\bar{n}^{\mu}$,
$W_{\eta}(x)=\bar{P}\Bigl{[}\exp\Bigl{(}-\mathrm{i}g\int_{0}^{\infty}\!\mathrm{d}s\,{\bar{n}}\cdot A(x+{\bar{n}}s)\Bigr{)}\Bigr{]}_{\eta}\,.$ (2.13)
The subscript $\eta$ indicates the presence of an additional rapidity regulator. In explicit calculations, we will use the so-called $\eta$ regulator of refs. [39, 40] whose action on a single gluon with momentum $\ell$ attached to the Wilson line amounts to inserting a factor of $\lvert\ell^{-}/\nu\rvert^{\eta}$ under phase-space or loop integrals over $\ell$, with $\nu$ a dimensionful rapidity scale. The final result for the bare matrix element depends on $\zeta/\nu^{2}$, with $\zeta\equiv(p^{-}/z)^{2}$ the Collins-Soper scale.
In terms of the bare collinear matrix element in eq. (2.2), the renormalized partonic heavy-quark TMD FF is given by
$d_{1\,Q/Q}(z,b_{T},\mu,\zeta)$ $=\lim_{\epsilon\to 0}Z_{\mathrm{UV}}(\mu,\zeta,\epsilon)\lim_{\eta\to 0}\Bigl{[}d_{1\,Q/Q}^{\mathrm{bare}}(z,b_{T},\epsilon,\eta,\zeta/\nu^{2})\,\sqrt{S}(b_{T},m,\epsilon,\eta,\nu)\Bigr{]}\,,$ (2.14)
where $S$ is the universal bare TMD soft function for the $\eta$ regulator, which cancels all poles of $\eta$ and the associated $\nu$ dependence within the second line and is independent of the heavy-quark mass up to secondary quark mass effects starting at two loops. By contrast, the $\overline{\text{MS}}$ renormalization factor $Z_{\mathrm{UV}}$ for quark TMDs is independent of the mass to all orders by RG consistency with the hard matching coefficient at the hard scattering energy $Q\sim p^{-}/z\gg m\sim k_{T}\sim 1/b_{T}$. While soft subtractions (known as zero bins in the SCET literature) generally need to be accounted for in both virtual and real collinear diagrams when computing the bare collinear matrix element itself, we have explicitly verified that they lead to scaleless integrals for our choice of regulator, also in the presence of the quark mass.
2.2 Nonvalence contributions
There are two further ways in which heavy quarks can participate in the TMD fragmentation process for $m\sim k_{T}\sim 1/b_{T}\gg\Lambda_{\mathrm{QCD}}$; these were sketched in ref. [1], but we spell them out here explicitly. In one case, the resolved final state is a heavy hadron $H$, while the parent parton $i=g,q,\bar{q}$ is light,
$D_{1\,H/i}(z_{H},b_{T},\mu,\zeta)=d_{1\,Q/i}(z_{H},b_{T},\mu,\zeta)\,\chi_{H}+\mathcal{O}(\Lambda_{\mathrm{QCD}})\,.$ (2.15)
Here $\chi_{H}$ is again the universal inclusive fragmentation probability for $Q\to H$, while the bare matching coefficient $d_{1\,Q/i}$ is obtained from eqs. (2.2) and (2.14) after replacing the heavy-quark fields in the bare correlator by suitable unpolarized combinations of gluon or light-quark fields.
In the other case, the resolved final-state hadron $h$ is light. In this case the TMD FF $D_{1\,h/i}$ for $i=g,q,\bar{Q},Q,\bar{Q}$ has to be matched onto collinear fragmentation functions $D_{h/j}$ at the scale $\Lambda_{\mathrm{QCD}}$ where all degrees of freedom $j=g,q,\bar{q}$ are light,
$D_{1\,h/i}(z_{H},b_{T},\mu,\zeta)=\frac{1}{z_{H}^{2}}\sum_{j}\int\!\frac{\mathrm{d}z}{z}\,\mathcal{J}_{j/i}(z,b_{T},m,\mu,\zeta)\,D_{h/j}\Bigl{(}\frac{z_{H}}{z},\mu\Bigr{)}+\mathcal{O}(\Lambda_{\mathrm{QCD}})\,.$ (2.16)
This has the standard form of matching TMD FFs onto collinear FFs, but note that we include the mass as an additional third argument of the matching coefficient to make explicit that even when $i$ is light, the heavy quark may in general contribute to the matching at $\mu\sim m\sim k_{T}$ through closed loops.
2.3 bHQET fragmentation factors and the large-mass limit
In ref. [1], we showed that for transverse momenta $k_{T}\ll m$, the TMD dynamics are governed by new (and in general nonperturbative) matrix elements defined in boosted HQET [41, 42] that we dubbed TMD fragmentation factors. Specifically, for $\Lambda_{\mathrm{QCD}}\lesssim k_{T}\ll m$ and counting
$1-z_{H}\sim 1\,,$ (2.17)
which amounts to integrating over wide bins in $z_{H}$ or taking low $z_{H}$ moments, the TMD FF for producing a heavy hadron $H$ off a heavy quark $Q$ factorizes as [1]
$D_{1\,H/Q}(z_{H},b_{T},\mu,\zeta)=\delta(1-z_{H})\,C_{m}(m,\mu,{\zeta})\,\chi_{1,H}\Bigl{(}b_{T},\mu,{\frac{\sqrt{\zeta}}{m}}\Bigr{)}+\mathcal{O}\Bigl{(}\frac{1}{m}\Bigr{)}\,,$ (2.18)
where $C_{m}(m,\mu,\zeta)$ is generated by separately matching collinear and soft modes at the scale $\mu\sim m$ onto HQET and QCD with $n_{\ell}$ flavors, respectively [43]; explicit expressions for $C_{m}$ and other perturbative ingredients to the order required for our perturbative checks are given in appendix B.
In eq. (2.18), the TMD dynamics are encoded in the unpolarized TMD fragmentation factor $\chi_{1,H}$. For $\Lambda_{\mathrm{QCD}}\ll k_{T}\sim 1/b_{T}$, we showed in ref. [1] that it is given by a product of the total fragmentation probability $\chi_{H}$ and a perturbative matching coefficient $C_{1}$,
$\chi_{1,H}(b_{T},\mu,{\rho})=C_{1}(b_{T},\mu,{\rho})\,\chi_{H}+\mathcal{O}(\Lambda_{\mathrm{QCD}}^{2}b_{T}^{2})\,,$ (2.19)
where $C_{1}$ will be the target of our calculation in section 4.3. Here we have also introduced the shorthand $\rho\equiv\bar{n}\cdot v=\sqrt{\zeta}/m$ for the boost of the hadron; in ref. [1] we argued from consistency with TMD factorization that $\chi_{1,H}$ develops an anomalous dependence on this variable governed by the Collins-Soper kernel, and we will make this expectation more explicit in section 4.1.
To evaluate $C_{1}$, we will use that – for fully general $\Lambda_{\mathrm{QCD}}\lesssim 1/b_{T}$ – the unpolarized TMD fragmentation factor $\chi_{1,H}$ reduces to a vacuum matrix element of Wilson lines when summed over all hadrons $H$ containing $Q$ [1]. Since $\sum_{H}\chi_{H}=1$ and $C_{1}$ in eq. (2.19) is independent of the hadronic state $H$, it is then straightforward to obtain $C_{1}$ by evaluating the Wilson line correlator at perturbative $b_{T}$.
At the partonic level, eqs. (2.18) and (2.19) together imply the following consistency condition for the partonic heavy-quark TMD FF in the limit $m\gg k_{T}\sim 1/b_{T}$ [1],
$d_{1\,Q/Q}(z,b_{T},\mu,\zeta)=\delta(1-z)\,C_{m}(m,\mu,{\zeta})\,C_{1}\Bigl{(}b_{T},\mu,{\frac{\sqrt{\zeta}}{m}}\Bigr{)}+\mathcal{O}\Bigl{(}\frac{1}{b_{T}m}\Bigr{)}\,,$ (2.20)
and in section 5.1 we will use this relation as a check of our one-loop results. By contrast, nonvalence (or disfavored) partonic heavy-quark TMD FFs become power-suppressed in the heavy-quark limit,
$i\neq Q\,:\quad d_{1\,Q/i}(z,b_{T},\mu,\zeta)=\mathcal{O}\Bigl{(}\frac{1}{b_{T}m}\Bigr{)}\,.$ (2.21)
The indirect effect ($D_{1\,h/i}$, $i\neq Q,\bar{Q}$) of the heavy quark on light-hadron production for $m\gg k_{T}$ is leading in $1/m$. It is governed by virtual contributions from collinear and soft so-called mass modes and thus follows exactly the TMD PDF case [44]. Finally, the direct contribution ($D_{1\,h/Q}$) to light-hadron production becomes strongly peaked at $z_{h}\to 0$ in the limit $m\gg k_{T}\gtrsim\Lambda_{\mathrm{QCD}}$, but for any finite value of $z_{h}$ (or argument $z$ of $\mathcal{J}_{j/Q}$ at the partonic level) is kinematically suppressed as $m\gg k_{T}$; we will verify this latter behavior in section 5.1, but leave a dedicated analysis of its $z_{h}\to 0$ behavior – which is known to be subtle in fragmentation [45] – to future work.
2.4 Consistency conditions in the light-quark limit
In the case when the heavy-quark is light compared to $k_{T}$, the matching at the scale $\mu\sim k_{T}$ takes exactly the standard form for matching TMD FFs onto twist-2 collinear FFs [1],
$D_{1\,\mathfrak{h}/Q}(z_{\mathfrak{h}},b_{T},\mu,\zeta)=\frac{1}{z_{\mathfrak{h}}^{2}}\sum_{i}\int\!\frac{\mathrm{d}z}{z}\,\mathcal{J}_{i/q}(z,b_{T},\mu,\zeta)\,D_{\mathfrak{h}/i}^{(n_{\ell}+1)}\Bigl{(}\frac{z_{\mathfrak{h}}}{z},\mu\Bigr{)}+\mathcal{O}(m^{2}b_{T}^{2})\,,$ (2.22)
where $\mathfrak{h}=h,H$ may be both a heavy or a light hadron and $D_{\mathfrak{h}/j}^{(n_{\ell}+1)}$ are collinear FFs in a theory with $n_{\ell}$ light and one massive flavor, i.e., the mass is the highest IR scale in the twist-2 matching. Importantly, the TMD FF matching coefficients $\mathcal{J}_{i/q}$ are independent of these details in the IR and are thus given by the universal TMD FF matching coefficients in a theory with $n_{f}=n_{\ell}+1$ light degrees of freedom, which are known to N3LO [46, 47]. The problem of the mass dependence is thus reduced to the well-understood behavior of the collinear FFs, which differs depending on whether $\mathfrak{h}$ is heavy or light [48, 49],
$D_{H/i}^{(n_{\ell}+1)}(z_{H},\mu)$ $=d_{Q/i}(z_{H},\mu)\,\chi_{H}+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{m}\Bigr{)}\,,$
$D_{h/i}^{(n_{\ell}+1)}(z_{h},\mu)$ $=\sum_{j}\int\!\frac{\mathrm{d}z}{z}\,\mathcal{M}_{j/i,T}(z,m,\mu)\,D_{h/j}^{(n)}\Bigl{(}\frac{z_{h}}{z},\mu\Bigr{)}\,.$ (2.23)
Here $d_{Q/i}(z_{H},\mu)$ is the partonic collinear heavy-quark FF and $\mathcal{M}_{j/i,T}$ is the timelike matching function governing the flavor decoupling in light-hadron collinear FFs.
Comparing eqs. (2.1), (2.15), and (2.16) to eqs. (2.22) and (2.4), we can read off the following leading-power behavior of the relevant perturbative heavy-quark TMD matrix elements in the limit $m\ll k_{T}\sim 1/b_{T}$ [1],
$d_{1\,Q/i}(z,b_{T},\mu,\zeta)$ $=\frac{1}{z^{2}}\sum_{j}\int\!\frac{\mathrm{d}z^{\prime}}{z^{\prime}}\,\mathcal{J}_{j/i}(z^{\prime},b_{T},\mu,\zeta)\,d_{Q/j}\Bigl{(}\frac{z}{z^{\prime}},\mu\Bigr{)}\,,$
$\mathcal{J}_{k/i}(z,b_{T},m,\mu,\zeta)$ $=\frac{1}{z^{2}}\sum_{j}\int\!\frac{\mathrm{d}z^{\prime}}{z^{\prime}}\,\mathcal{J}_{j/i}(z^{\prime},b_{T},\mu,\zeta)\,\mathcal{M}_{k/j,T}\Bigl{(}\frac{z}{z^{\prime}},m,\mu\Bigr{)}\,.$ (2.24)
In section 5.2, we will use these expressions, together with the various known ingredients on the right-hand side, to perform cross checks of our one-loop results in section 3 in all channels.
3 Partonic heavy-quark TMD fragmentation at NLO
For ease of calculation and to make contact with the splitting probability in eq. (1.1), we will perform our calculation of the bare collinear matrix element as well as its renormalization in momentum space, i.e., as a function of $z$ and $k_{T}$, and later compute various integral transforms of the renormalized object directly. Passing to momentum space, the one-loop correction to the renormalized heavy-quark TMD FF in eq. (2.14) reads
$d_{1\,Q/Q}^{(1)}(z,k_{T},\mu,\zeta)$ $=\lim_{\epsilon\to 0}\biggl{\{}\delta(1-z)\,\frac{1}{\pi}\delta(k_{T}^{2})\,Z_{\mathrm{UV}}^{(1)}(\mu,\zeta,\epsilon)$ (3.1)
$\quad+\lim_{\eta\to 0}\Bigl{[}d_{1\,Q/Q}^{\mathrm{bare}\,(1)}(z,k_{T},\epsilon,\eta,\zeta/\nu^{2})+\delta(1-z)\,\tfrac{1}{2}S^{(1)}(k_{T},\epsilon,\eta,\nu)\Bigr{]}\biggr{\}}\,,$
where we have used that $d_{1\,Q/Q}^{(0)}=\delta(1-z)\,\frac{1}{\pi}\delta(k_{T}^{2})$ at tree level. (Here and in the following we use the same symbol for functions of $b_{T}$ and their Fourier transforms, as the meaning is always clear from the argument.) Note that in our implementation of dimensional regularization in the transverse plane, we make use of the azimuthal symmetry of the unpolarized TMD FF and take the curly braces in eq. (3.1) to be a density in $\pi k_{T}^{2}$ for convenience, i.e., they have an integer mass dimension of $-2$, where the factor of $\pi$ ensures that final results in $d=4$ are properly normalized azimuthally symmetric densities in vectorial $\vec{k}_{T}$ [50, 51]. With these preliminaries, the bare collinear matrix element in momentum space is given by
[…]
×$\quad\times\big{\langle}QX\big{\lvert}\bigl{[}\delta\bigl{(}p^{-}/z+\mathrm{i}\partial^{-}\bigr{)}\,\delta^{(2-2\epsilon)}\bigl{(}k_{\perp}^{\mu}+\mathrm{i}\partial_{\perp}^{\mu}\bigr{)}\,\bar{\psi}_{Q}\,W_{\eta}\bigr{]}\big{\rvert}0\big{\rangle}\Bigr{]}\,.$
Here $\mathrm{d}\Omega_{2-2\epsilon}$ is the solid angle element for $k_{\perp}^{\mu}$ in $2-2\epsilon$ dimensions and the $\delta$ functions involving derivative operators on the second line act on the fields to their right, i.e., they fix the total minus and perpendicular momentum injected into the compound operator.
Compared to a direct calculation in position space, which involves expanding Bessel-like hypergeometric functions in the dimensional regulator, the main challenge in the momentum-space calculation is the careful distributional treatment of the singularity structure prior to expanding in $\epsilon$, which we address in section 3.1. Renormalized results in momentum space are presented in section 3.2, which we independently verify and further simplify by moving to cumulant space in section 3.3. Results in $b_{T}$ space are then obtained by a straightforward integral transform mapping finite functions onto functions, and are given in section 3.4.
3.1 Bare real-emission diagrams and distributional expansion
[…]
The two contributing real-emission diagrams in Feynman gauge are given in figure -1078 (a) and (b), where $X=g$ is a single gluon. Including the mirror diagram for (a), their contributions to $d_{1\,Q/Q}^{\mathrm{bare}\,(1)}$ evaluate to
$d_{1\,Q/Q}^{\mathrm{bare}\,(a)}$ $=\frac{\alpha_{s}C_{F}}{4\pi}\frac{e^{\epsilon\gamma_{E}}}{\Gamma(1-\epsilon)}\Bigl{(}\frac{\mu}{k_{T}}\Bigr{)}^{2\epsilon}\Bigl{(}\frac{\sqrt{\zeta}}{\nu}\Bigr{)}^{-\eta}\frac{1}{\pi z^{2-2\epsilon}}\frac{z^{\eta}}{\bar{z}^{1+\eta}}\frac{4z^{3}}{k_{T}^{2}z^{2}+m^{2}\bar{z}^{2}}\,,$ (3.11)
[…]
×$\quad\times\Bigl{[}\frac{k_{T}^{2}z^{2}+m^{2}(1-4z+z^{2})-\epsilon(k_{T}^{2}z^{2}+m^{2}\bar{z}^{2})}{(k_{T}^{2}z^{2}+m^{2}\bar{z}^{2})^{2}}\Bigr{]}\,,$ (3.12)
where we defined the shorthand $\bar{z}\equiv 1-z$ and $\alpha_{s}=\alpha_{s}(\mu)$ is the renormalized $\overline{\text{MS}}$ coupling. Note that to our working order, all results in this paper are independent of whether they are expressed in terms of $\alpha_{s}^{(n_{f})}$ or $\alpha_{s}^{(n_{\ell})}$. We suppress overall factors of $\Theta(z)$, $\Theta(\bar{z})$, or $\Theta(k_{T}^{2})$ from the final-state on-shell conditions in the following, but stress that they are formally present in all the distributional identities we use or derive, and also multiply all finite remainder terms.
The two seemingly simple expressions above feature an intricate interplay of singularities as $k_{T}\to 0$ and/or $z\to 1$ ($\bar{z}\to 0$), which in particular arise as the quark propagator goes on shell for $k_{T}^{2}z^{2}+m^{2}\bar{z}^{2}\to 0$. At this step, the singularities are regulated by powers of $\bar{z}^{-\eta}$, dimensional regularization, or both. However, if we restrict to finite $z<1$ and $k_{T}>0$, the limits $\eta\to 0$ and $\epsilon\to 0$ can be taken right away and all other contributions in the renormalized one-loop formula including the virtual diagrams drop out in eq. (3.1), leaving behind a finite result from the sum of (a) and (b),
$z<1~{}\text{and}~{}k_{T}>0\,:\quad d_{1\,Q/Q}^{(1)}(z,k_{T},\mu,\zeta)=\frac{\alpha_{s}C_{F}}{4\pi}\frac{2}{\pi}\frac{k_{T}^{2}z^{2}(1+z^{2})+m^{2}\bar{z}^{4}}{\bar{z}(k_{T}^{2}z^{2}+m^{2}\bar{z}^{2})^{2}}\,.$ (3.13)
This is, of course, exactly the differential splitting probability for $Q\bar{Q}\to(Qg)\bar{Q}$ in the quasi-collinear limit that we considered in eq. (1.1), but now the field-theoretic definition of $d_{1\,Q/Q}$ provides us with explicit regulators controlling the singular limits. It is in this sense that the heavy-quark TMD FF, after including the virtual diagram and performing the renormalization, will provide a fully differential extension of the splitting probability into the unresolved limit(s) that is finite in $d=4$, universal, and embedded in factorization theorems.
To proceed, we first expand the Wilson line diagram (a) in the rapidity regulator $\eta$, using the standard identity
$\frac{z^{\eta}}{\bar{z}^{1+\eta}}=-\frac{\delta(\bar{z})}{\eta}+\mathcal{L}_{0}(\bar{z})+\mathcal{O}(\eta)\,.$ (3.14)
This expansion is valid here because the remaining terms in eq. (3.11) are still dimensionally regulated at $\bar{z}=0$ as $k_{T}\to 0$ i.e., the remainder of the expression is a good test function for general $\epsilon$. Since the mass enters diagram (a) only through the combination $m\bar{z}$, the action of the $\delta(\bar{z})/\eta$ term on the rest of the diagram is to reduce it to the massless limit. This in particular means that the coefficient of the $1/\eta$ pole (including its exact $\epsilon$ dependence) is equal to the one for the massless TMD FF, as it must for it to cancel the pole in the $\eta$-regulated soft function, which is independent of the mass at this order. For our implementation of dimensional regularization, the soft function is given by
$\frac{1}{2}S^{(1)}(k_{T},\epsilon,\eta,\nu)=\frac{\alpha_{s}C_{F}}{4\pi}\frac{e^{\epsilon\gamma_{E}}}{\Gamma(1-\epsilon)}\Bigl{(}\frac{\mu}{k_{T}}\Bigr{)}^{2\epsilon}\frac{1}{\pi k_{T}^{2}}\Bigl{[}+\frac{4}{\eta}+2\ln\frac{\nu^{2}}{k_{T}^{2}}+\mathcal{O}(\eta)\Bigr{]}\,.$ (3.15)
Combining this with diagram (a), the $\eta\to 0$ limit can be taken, and we find
$d_{1\,Q/Q}^{\mathrm{bare}\,(a+s)}$ $\equiv\lim_{\eta\to 0}\Bigl{[}d_{1\,Q/Q}^{\mathrm{bare}\,(a)}(z,k_{T},\epsilon,\eta,\zeta/\nu^{2})+\delta(1-z)\,\tfrac{1}{2}S^{(1)}(k_{T},\epsilon,\eta,\nu)\Bigr{]}$ (3.16)
$=\frac{\alpha_{s}C_{F}}{4\pi}\frac{e^{\epsilon\gamma_{E}}}{\Gamma(1-\epsilon)}\frac{1}{\pi z^{2-2\epsilon}}\biggl{\{}\frac{\mu^{2\epsilon}}{k_{T}^{2+2\epsilon}}\Bigl{[}2\ln\frac{\zeta}{k_{T}^{2}}\delta(\bar{z})+4\mathcal{L}_{0}(\bar{z})\Bigr{]}+\frac{\mu^{2\epsilon}}{m^{2+2\epsilon}}\,f^{(a)}(k_{T}^{2}/m^{2},z,\epsilon)\biggr{\}}\,.$
Here the first term in curly braces, which is left behind from the cancellation against the soft function and the same as for the massless case, is straightforward to expand in $\epsilon$ using identities like eqs. (A.4) and (A.5). This produces, among others, a term $\tfrac{1}{\epsilon^{2}}\delta(\bar{z})\delta(k_{T}^{2})$ as well as a term $\tfrac{1}{\epsilon}\mathcal{L}_{0}(\bar{z})\delta(k_{T}^{2})$. The nontrivial part is to expand the dimensionless function
$f^{(a)}(x,z,\epsilon)\equiv x^{-1-\epsilon}\frac{-4+4z-4xz^{2}}{\bar{z}^{2}+xz^{2}}$ (3.17)
in a Laurent series in $\epsilon$ involving plus distributions of $x$ and $z$. To do so, we recall that plus distributions in one variable $x$ formally arise in perturbative calculations as
$f(x,\epsilon)=\bigl{[}f(x,\epsilon)\bigr{]}_{+}+\delta(x)\,F(\epsilon)\,,\qquad F(\epsilon)=\int_{0}^{1}\!\mathrm{d}x^{\prime}\,f(x^{\prime},\epsilon)\,,$ (3.18)
where $f(x,\epsilon)\sim x^{-1-\epsilon}+\mathcal{O}(x^{0})$ has a dimensionally regulated singular limit $x\to 0$ on the first left-hand side. By contrast, on the right-hand side $f(x,\epsilon)$ may be expanded in $\epsilon$ since each term in the expansion is now dimensionally regulated, whereas the explicit poles are typically isolated in the Laurent series of the total integral $F(\epsilon)$. Of course, this is precisely how relations like eq. (3.14) are derived.
The identity in eq. (3.18) generalizes to our case of two variables $x$ and $z$, cf. ref. [52],
$f(x,z,\epsilon)=\bigl{[}f(x,z,\epsilon)\bigr{]}_{+,+}+\delta(x)\bigl{[}F_{x}(z,\epsilon)\bigr{]}_{+}+\delta(\bar{z})\bigl{[}F_{z}(x,\epsilon)\bigr{]}_{+}+\delta(x)\,\delta(\bar{z})\,F_{xz}(\epsilon)\,,$ (3.19)
where
$F_{x}(z,\epsilon)$ $\equiv\int_{0}^{1}\!\mathrm{d}z^{\prime}\,f(x,z^{\prime},\epsilon)\,,\qquad F_{z}(x,\epsilon)\equiv\int_{0}^{1}\!\mathrm{d}x^{\prime}\,f(x^{\prime},z,\epsilon)\,,$
$\qquad F_{xz}(\epsilon)\equiv\int_{0}^{1}\!\mathrm{d}x\,\int_{0}^{1}\!\mathrm{d}z^{\prime}\,f(x,z^{\prime},\epsilon)$ (3.20)
and the two-dimensional plus bracket $[\dots]_{+,+}$ is defined through its action on test functions $g(x,z)$ as
$\int_{0}^{1}\!\mathrm{d}x\int_{0}^{1}\!\mathrm{d}z\,\bigl{[}f(x,z,\epsilon)\bigr{]}_{+,+}\,g(x,z)$
$\equiv\int_{0}^{1}\!\mathrm{d}x\int_{0}^{1}\!\mathrm{d}z\,f(x,z,\epsilon)\,\bigl{[}g(x,z)-g(0,z)-g(x,0)+g(0,0)\bigr{]}\,.$ (3.21)
Note that in our case with $x=k_{T}^{2}/m^{2}$, the integration domain can also extend to $x_{*}>x>1$, in which case the $[\dots]_{+,+}$ bracket reduces to a one-dimensional one,
$\int_{1}^{x_{*}}\!\mathrm{d}x\int_{0}^{1}\!\mathrm{d}z\,\bigl{[}f(x,z,\epsilon)\bigr{]}_{+,+}\,g(x,z)\equiv\int_{1}^{x_{*}}\!\mathrm{d}x\int_{0}^{1}\!\mathrm{d}z\,f(x,z,\epsilon)\,\bigl{[}g(x,z)-g(x,0)\bigr{]}\,,$ (3.22)
while integration domains with other shapes can be implemented through suitable test functions. So far, eq. (3.19) together with eq. (3.1) is a tautology; however, as in the one-dimensional case in eq. (3.18), the crucial step is that the expansion in $\epsilon$ may now be performed under each of the various plus brackets.
For the case of $f=f^{(a)}$ as defined in eq. (3.17), the various integrals in eq. (3.1) are all straightforward to evaluate for general $\epsilon$ in terms of incomplete Beta and digamma functions. If we restrict to their expansions in $\epsilon$ for brevity, eq. (3.19) evaluates to
$f^{(a)}(x,z,\epsilon)$ $=\bigl{[}f^{(a)}(x,z,0)\bigr{]}_{+,+}+\delta(x)\biggl{\{}\frac{4}{\epsilon}\mathcal{L}_{0}(\bar{z})+4\biggl{[}\frac{4z\ln\bigl{(}1+z^{2}/\bar{z}^{2}\bigr{)}}{\bar{z}}\biggr{]}_{+}\biggr{\}}$
$\quad+\delta(\bar{z})\biggl{[}\frac{2(1-x)\ln x-4(\pi\sqrt{x}+x+x^{2})}{x(1+x)^{2}}\biggr{]}_{+}$
$\quad+\delta(\bar{z})\delta(x)\biggl{\{}-\frac{2}{\epsilon^{2}}-2\pi-\frac{5\pi^{2}}{6}\biggr{\}}+\mathcal{O}(\epsilon)\,.$ (3.23)
In this form, the double pole $1/\epsilon^{2}$ is isolated at $\bar{z}=x=0$. The term proportional to $\tfrac{1}{\epsilon}\mathcal{L}_{0}(\bar{z})$ along the $x=0$ boundary ends up canceling a corresponding term from the $\epsilon$ expansion of the first term in curly braces on the second line of eq. (3.16), such that $d_{1\,Q/Q}^{\mathrm{bare}\,(a+s)}$ is finite as $\epsilon\to 0$ for $z<1$ for any $x\geq 0$. The double pole $\tfrac{1}{\epsilon}\delta(\bar{z})\delta(k_{T}^{2})$ similarly cancels.
The second diagram in figure -1078 does not require rapidity regularization. In order to expand it in the dimensional regulator, we repeat the above procedure, where we define
$d_{1\,Q/Q}^{\mathrm{bare}\,(b)}$ $\equiv\frac{\alpha_{s}C_{F}}{4\pi}\frac{e^{\epsilon\gamma_{E}}}{\Gamma(1-\epsilon)}\frac{1}{\pi z^{2-2\epsilon}}\frac{\mu^{2\epsilon}}{m^{2+2\epsilon}}\,f^{(b)}(k_{T}^{2}/m^{2},z,\epsilon)\,,$
$f^{(b)}(x,z,\epsilon)$ $=2z^{2}\bar{z}\,x^{-2\epsilon}\frac{xz^{2}+(1-4z+z^{2})-\epsilon(xz^{2}+\bar{z}^{2})}{(xz^{2}+\bar{z}^{2})^{2}}\,.$ (3.24)
The integrals in eq. (3.1) for $f=f^{(b)}$ can again be performed including their exact dependence on $\epsilon$, leading to digamma and ${}_{2}F_{1}$ hypergeometric functions. Expanding the latter in $\epsilon$ using the HypExp 2.0 package [53], we find the following distributional expansion of the dimensionless function $f^{(b)}$,
$f^{(b)}(x,z,\epsilon)$ $=\bigl{[}f^{(b)}(x,z,0)\bigr{]}_{+,+}+\delta(x)\biggl{[}\frac{4z^{3}}{\bar{z}(2\bar{z}z-1)}+2\bar{z}\ln\Bigl{(}1+\frac{z^{2}}{\bar{z}^{2}}\Bigr{)}\biggr{]}_{+}$ (3.25)
$\quad+\delta(\bar{z})\biggl{[}\frac{x(5+x)\ln x+(-1+\pi\sqrt{x}-x)(2-3x-3x^{2})}{x(1+x)^{3}}\biggr{]}_{+}$ (3.26)
$\quad+\delta(\bar{z})\delta(x)\biggl{\{}\frac{2}{\epsilon}+1+\frac{3\pi}{2}\biggr{\}}+\mathcal{O}(\epsilon)\,.$ (3.27)
Combining eqs. (3.1) and (3.25) produces an explicit result for the sum of real-emission diagrams expanded in $\epsilon$. The finite terms involving the one and two-dimensional plus distributions are lengthy and reappear in our renormalized result below, so we only quote the poles for reference:
$d_{1\,Q/Q}^{\mathrm{bare}\,(a+s)}+d_{1\,Q/Q}^{\mathrm{bare}\,(b)}=\frac{\alpha_{s}C_{F}}{4\pi}\frac{1}{\pi z^{2}}\,\biggl{\{}\delta(\bar{z})\delta(k_{T}^{2})\Bigl{(}-\frac{2}{\epsilon}\ln\frac{\zeta}{m^{2}}+\frac{2}{\epsilon}\Bigr{)}+\mathcal{O}(\epsilon^{0})\biggr{\}}\,.$ (3.28)
3.2 Virtual contributions and UV renormalization
The virtual diagram (c) from figure -1078 is common to any collinear matrix element involving massive fermions and can be evaluated in a straightforward way using integration by contours for the $\ell^{+}$ component of the loop momentum $\ell$, Feynman parametrization for the $\ell_{\perp}$ integral, and an analytic $\ell^{-}$ integral at the end. Including the mirror diagram, the result reads
$d_{1\,Q/Q}^{\mathrm{bare}\,(c)}=\frac{\alpha_{s}C_{F}}{4\pi}\,\delta(1-z)\frac{1}{\pi}\delta(k_{T}^{2})\,\frac{2e^{\epsilon\gamma_{E}}\Gamma(\epsilon)}{\epsilon(1-2\epsilon)}\frac{\mu^{2\epsilon}}{m^{2\epsilon}}\,,$ (3.29)
and contributes a double $1/\epsilon^{2}$ pole. So far our calculation was performed in terms of bare quark fields, and we still have to account for eq. (2.12). Using the standard on-shell renormalization factor for a quark field with pole mass $m$ in dimensional regularization,
$Z_{\psi,\mathrm{OS}}\equiv 1+Z_{\psi,\mathrm{OS}}^{(1)}+\mathcal{O}(\alpha_{s}^{2})\,,\qquad Z_{\psi,\mathrm{OS}}^{(1)}=-\,\frac{\alpha_{s}C_{F}}{4\pi}\Big{(}\frac{3}{\epsilon}+3\ln\frac{\mu^{2}}{m^{2}}+4\Big{)}$ (3.30)
the rapidity-renormalized collinear correlator on the second line of eq. (3.1) is given by
$\lim_{\eta\to 0}\Bigl{[}d_{1\,Q/Q}^{\mathrm{bare}\,(1)}(z,k_{T},\epsilon,\eta,\zeta/\nu^{2})+\delta(1-z)\,\tfrac{1}{2}S^{(1)}(k_{T},\epsilon,\eta,\nu)\Bigr{]}$
$=d_{1\,Q/Q}^{\mathrm{bare}\,(a+s)}+d_{1\,Q/Q}^{\mathrm{bare}\,(b)}+d_{1\,Q/Q}^{\mathrm{bare}\,(c)}+\delta(1-z)\frac{1}{\pi}\delta(k_{T}^{2})Z_{\psi}^{(1)}$
$=\frac{\alpha_{s}C_{F}}{4\pi}\frac{1}{\pi z^{2}}\,\biggl{\{}\delta(\bar{z})\delta(k_{T}^{2})\Bigl{(}\frac{2}{\epsilon^{2}}+\frac{3}{\epsilon}-\frac{2}{\epsilon}\ln\frac{\zeta}{\mu^{2}}\Bigr{)}+\mathcal{O}(\epsilon^{0})\biggr{\}}\,,$ (3.31)
and we continue to only quote the poles for brevity. The $\overline{\text{MS}}$ UV renormalization factor for quark TMD PDFs and FFs is likewise well known and independent of the mass [54],
$Z_{\mathrm{UV}}^{(1)}(\mu,\zeta,\epsilon)=-\frac{\alpha_{s}C_{F}}{4\pi}\biggl{\{}\frac{2}{\epsilon^{2}}+\frac{3}{\epsilon}-\frac{2}{\epsilon}\ln\frac{\zeta}{\mu^{2}}\biggr{\}}\,.$ (3.32)
It precisely cancels the poles from eq. (3.2) when inserting both into eq. (3.1), as expected. We thus confirm by an explicit one-loop calculation that the TMD FF for heavy quarks obeys standard TMD evolution equations, with secondary mass effects in the Collins-Soper kernel at higher loop orders understood [44].
Taking the $\epsilon\to 0$ limit, we obtain our final result for the $\mathcal{O}(\alpha_{s})$ correction to the renormalized partonic heavy-quark TMD FF:
$d_{1\,Q/Q}^{(1)}(z,k_{T},\mu,\zeta)$
$=\frac{\alpha_{s}C_{F}}{4\pi}\frac{1}{\pi z^{2}}\,\biggl{\{}\delta(\bar{z})\Bigl{[}2\ln\frac{\zeta}{\mu^{2}}\mathcal{L}_{0}(k_{T}^{2},\mu^{2})+2\ln\frac{\mu^{2}}{m^{2}}\mathcal{L}_{0}(k_{T}^{2},m^{2})-\ln^{2}\frac{\mu^{2}}{m^{2}}\delta(k_{T}^{2})+3\ln\frac{\mu^{2}}{m^{2}}\delta(k_{T}^{2})\Bigr{]}$
$\qquad+\frac{1}{m^{2}}\biggl{[}\frac{2xz^{4}(1+z^{2})+2z^{2}\bar{z}^{4}}{\bar{z}(xz^{2}+\bar{z}^{2})^{2}}\biggr{]}_{+,+}+\delta(k_{T}^{2})\biggl{[}\frac{2(1+z^{2})}{\bar{z}}\ln\Bigl{(}1+\frac{z^{2}}{\bar{z}^{2}}\Bigr{)}-\frac{4z^{3}}{\bar{z}(1-2\bar{z}z)}\biggr{]}_{+}$
$\qquad+\delta(\bar{z})\,\frac{1}{m^{2}}\biggl{[}-\frac{2+3x+4x^{2}+3x^{3}+\pi\sqrt{x}(2+7x+x^{2})+x(1+7x+2x^{2})\ln x}{x(1+x)^{3}}\biggr{]}_{+}$
$\qquad+\delta(\bar{z})\,\delta(k_{T}^{2})\Bigl{(}5-\frac{\pi}{2}-\frac{2\pi^{2}}{3}\Bigr{)}\biggr{\}}\,.$ (3.33)
Here we still use the dimensionless variable $x\equiv k_{T}^{2}/m^{2}$. To simplify the result we have combined terms under +,++,++ , + and $+$ brackets as much as possible, exploiting their linearity, reexpressed plus distributions $\mathcal{L}_{n}(k_{T}^{2},\mu^{2})$ in terms of $\mathcal{L}_{n}(k_{T}^{2},m^{2})$ and logarithms of $\mu^{2}/m^{2}$, and used that products of one-dimensional plus distributions can be written as (trivial) two-dimensional ones, i.e.,
$\mathcal{L}_{0}(\bar{z})\,\mathcal{L}_{0}(k_{T}^{2},m^{2})=\mathcal{L}_{0}(\bar{z})\,\frac{1}{m^{2}}\mathcal{L}_{0}(k_{T}^{2}/m^{2})=\frac{1}{m^{2}}\Bigl{[}\frac{1}{x}\frac{1}{\bar{z}}\Bigr{]}_{+,+}\,.$ (3.34)
Eq. (3.2) is the main result of this section. Up to the terms on the first line of the right-hand side, which are predicted by the RGE and vanish for $\zeta=\mu^{2}=m^{2}$, we have cast the heavy-quark TMD FF precisely in the form of eq. (3.19), i.e., as a two-dimensional plus distribution with nontrivial functions of $z$ and $k_{T}$ living on the respective boundaries and an overall boundary contribution at $z=1$ and $k_{T}=0$. This structure was already illustrated in figure -1079, where the partial integrals over one of the variables can immediately be read off from the respective boundary terms. Of course, the total contribution in the bulk at $z<1$ and $k_{T}>0$, which can simply be read off from the content of the $[\dots]_{+,+}$ brackets accounting for the prefactor of $1/z^{2}$, is still equal to the simple result we found in eq. (3.13) and thus to the quasi-collinear splitting probability in eq. (1.1).
3.3 Results in cumulant space
Fixed-order subtraction methods, as well as certain $q_{T}$ resummation formalisms like the RadISH approach [55], require the cumulative distribution in $k_{T}$ integrated over transverse momenta $\vec{k}_{T}$ with $\lvert\vec{k}_{T}\rvert\leq{k_{T}^{\mathrm{cut}}}$ as an input,
$d_{1\,Q/Q}(z,{k_{T}^{\mathrm{cut}}},\mu,\zeta)\equiv\pi\!\int^{({k_{T}^{\mathrm{cut}}})^{2}}\!\mathrm{d}(k_{T}^{2})\,d_{1\,Q/Q}(z,k_{T},\mu,\zeta)$ (3.35)
often simply referred to as the cumulant. As for the Fourier transform we will indicate the cumulative distribution simply through its argument ${k_{T}^{\mathrm{cut}}}$.
The organization of our final result in eq. (3.2) in terms of plus distributions defined on rectangular domains in $(z,k_{T}^{2})$ makes it very straightforward to compute the cumulative distribution because the cumulant constitutes a special case of eq. (3.1) for a unit test function. Specifically, for plus-bracketed functions $f$ of either a single variable $x=k_{T}^{2}/m^{2}$ or of both $x$ and $z$, we have the following integrals up to $x_{\mathrm{cut}}\equiv({k_{T}^{\mathrm{cut}}}/m)^{2}$:
$\int^{x_{\mathrm{cut}}}\!\mathrm{d}x\,\bigl{[}f(x,z)\bigr{]}_{+,+}$ $=\biggl{[}\int_{1}^{x_{\mathrm{cut}}}\!f(x,z)\biggr{]}_{+}\,,$
$\int^{x_{\mathrm{cut}}}\!\mathrm{d}x\,\bigl{[}f(x)\bigr{]}_{+}$ $=\int_{1}^{x_{\mathrm{cut}}}\!f(x)\,,$ (3.36)
where the remaining one-dimensional plus bracket on the right-hand side of the first line refers to the $z$ dependence of the integral inside it, while the latter’s $x_{\mathrm{cut}}$ dependence is function valued.
The integrals in eq. (3.3) are straightforward to evaluate for the respective terms in eq. (3.2). For purposes of our presentation, we find it convenient to further decompose the $z$ plus distribution arising from the cumulant of the $[\dots]_{+,+}$ term as
$\bigl{[}s(x_{\mathrm{cut}},z)+r(x_{\mathrm{cut}},z)\bigr{]}_{+}=\bigl{[}s(x_{\mathrm{cut}},z)\bigr{]}+r(x_{\mathrm{cut}},z)-\delta(\bar{z})\int_{0}^{1}\!\mathrm{d}z^{\prime}\,r(x_{\mathrm{cut}},z^{\prime})\,,$ (3.37)
where $s(x_{\mathrm{cut}},z)$ is defined as the leading term under the brackets for $\bar{z}\ll x_{\mathrm{cut}}\sim 1$, with homogeneous $\sim 1/\bar{z}$ scaling, while $r(x_{\mathrm{cut}},z)$ is defined as the remainder that is at most logarithmically divergent as $z\to 1$. Notably, we find that in this operation, the $x_{\mathrm{cut}}$ dependence of $s(z)$ is purely logarithmic, while the power-like $x_{\mathrm{cut}}$ dependence of the integrated $r(x_{\mathrm{cut}},z^{\prime})$ multiplying $\delta(\bar{z})$ precisely compensates that of the complicated cumulant integral of the one-dimensional plus distribution in $x$ in eq. (3.2).
Collecting terms, our final result for the NLO correction to the renormalized heavy-quark TMD FF in cumulant space reads
$d_{1\,Q/Q}^{\,(1)}(z,{k_{T}^{\mathrm{cut}}},\mu,\zeta)$
$=\frac{\alpha_{s}C_{F}}{4\pi}\frac{1}{z^{2}}\bigg{\{}\delta(1-z)\bigg{[}-4\ln\frac{\mu}{{k_{T}^{\mathrm{cut}}}}\ln\frac{\zeta}{m^{2}}+4\ln^{2}\frac{\mu}{m}+6\ln\frac{\mu}{m}-4\ln^{2}\frac{{k_{T}^{\mathrm{cut}}}}{m}+4-\frac{\pi^{2}}{6}\bigg{]}$
$\qquad\qquad\qquad-4\Bigl{(}1-2\ln\frac{{k_{T}^{\mathrm{cut}}}}{m}\Bigr{)}\,\mathcal{L}_{0}(1-z)-8\mathcal{L}_{1}(1-z)+\mathcal{R}\Bigl{[}z,\Bigl{(}\frac{{k_{T}^{\mathrm{cut}}}}{m}\Bigr{)}^{\!2\,}\Bigr{]}\bigg{\}}\,,$ (3.38)
where we defined
$\mathcal{R}(z,x_{\mathrm{cut}})\equiv$ $\,\,\frac{4-4z(1-x_{\mathrm{cut}}z)}{\bar{z}^{2}+x_{\mathrm{cut}}z^{2}}+4\ln\frac{\bar{z}^{2}}{x_{\mathrm{cut}}}+\frac{1}{\bar{z}}\Bigl{[}2(1+z^{2})\ln\Bigl{(}1+\frac{x_{\mathrm{cut}}z^{2}}{\bar{z}^{2}}\Bigr{)}-4z\ln\frac{x_{\mathrm{cut}}}{\bar{z}^{2}}\Bigr{]}\,.$ (3.39)
We note that $\mathcal{R}(z,x_{\mathrm{cut}})=\mathcal{O}(\bar{z}^{0})$ for $z\to 1$ and also $\mathcal{R}(z,x_{\mathrm{cut}})=\mathcal{O}(z^{0})$ for $z\to 0$, so it is integrable in all limits. For later reference we note its total integral,
$\int_{0}^{1}\!\mathrm{d}z\,\mathcal{R}(z,x_{\mathrm{cut}})$ $=-4\mathrm{Li}_{2}\Bigl{(}1+\frac{1}{\mathrm{i}\sqrt{x_{\mathrm{cut}}}}\Bigr{)}-4\mathrm{Li}_{2}\Bigl{(}1-\frac{1}{\mathrm{i}\sqrt{x_{\mathrm{cut}}}}\Bigr{)}$
$\quad-\frac{3x_{\mathrm{cut}}^{2}+3x_{\mathrm{cut}}+2}{(1+x_{\mathrm{cut}})^{2}}\ln x_{\mathrm{cut}}+\frac{2x_{\mathrm{cut}}(x_{\mathrm{cut}}-\pi\sqrt{x_{\mathrm{cut}}}+1)}{(1+x_{\mathrm{cut}})^{2}}\,.$ (3.40)
To check eq. (3.3), we have performed an independent calculation of the bare collinear correlator in dimensional regularization and renormalized it directly in cumulant space, This approach involves more complicated hypergeometric functions of combinations of $x_{\mathrm{cut}}$ and $z$ in intermediate steps; in particular, the challenge in this case is that their expansion in $\epsilon$ produces terms that are singular as $z\to 1$, even after subtracting terms that scale as $1/\bar{z}^{1+\epsilon}$ for general $\epsilon$. This means that the full hypergeometric result in dimensional regularization needs to be treated as in eq. (3.18) before expanding in $\epsilon$. The final result is in full agreement with eq. (3.3).
Eq. (3.3) has the – at first sight – remarkable property that the dependence of the coefficient of $\delta(1-z)$ on ${k_{T}^{\mathrm{cut}}}/m$ is purely logarithmic, while a power-like dependence on ${k_{T}^{\mathrm{cut}}}/m$ would in general be allowed by our power counting assumption ${k_{T}^{\mathrm{cut}}}\sim m$ for the heavy-quark TMD FF. Instead, after the cancellations discussed below eq. (3.37) have taken place, the nontrivial power-like dependence on ${k_{T}^{\mathrm{cut}}}/m$ arises solely in the $\mathcal{R}(z,x_{\mathrm{cut}})$ term, which is regular as $z\to 1$. In section 6 we will explain this fact from the perspective of joint threshold and $k_{T}$ factorization, and give all-order expressions for the $z\to 1$ limit of $d_{1\,Q/Q}$ at fixed $k_{T}/m\sim b_{T}m\sim 1$.
3.4 Results in position space
We now move on to position-space results in terms of the $\vec{b}_{T}$ variable Fourier-conjugate to $\vec{k}_{T}$, which are the key input for producing resummed predictions from solving multiplicative $b_{T}$-space RGEs and Fourier-transforming the final result. As is standard, the two-dimensional Fourier transform of the azimuthally symmetric renormalized TMD FF at hand is simply given by a Fourier-Bessel (or Hankel) transform of zeroth order,
$d_{1\,Q/Q}(z,b_{T},\mu,\zeta)=2\pi\int_{0}^{\infty}\!\mathrm{d}k_{T}\,k_{T}\,J_{0}(b_{T}k_{T})\,d_{1\,Q/Q}(z,k_{T},\mu,\zeta)\,,$ (3.41)
where $J_{n}$ is the $n$th-order Bessel function of the first kind. It is convenient to integrate the above relation by parts,
$d_{1\,Q/Q}(z,b_{T},\mu,\zeta)=b_{T}\int_{0}^{\infty}\!\mathrm{d}{k_{T}^{\mathrm{cut}}}\,J_{1}(b_{T}{k_{T}^{\mathrm{cut}}})\,d_{1\,Q/Q}(z,{k_{T}^{\mathrm{cut}}},\mu,\zeta)\,.$ (3.42)
This lets us compute the Fourier-space result directly from the cumulant in eq. (3.3) obtained in the previous section, which already had function-valued (and simpler) ${k_{T}^{\mathrm{cut}}}$ dependence, cf. the more complicated original momentum-space result in eq. (3.2).
Evaluating the relevant Hankel transforms, our result for the NLO correction to the renormalized heavy-quark TMD FF in position space reads
$d_{1\,Q/Q}^{\,(1)}(z,b_{T},\mu,\zeta)$ $=\frac{\alpha_{s}C_{F}}{4\pi}\frac{1}{z^{2}}\bigg{\{}\delta(1-z)\bigg{[}-2L_{b}\ln\frac{\zeta}{m^{2}}+4\ln^{2}\frac{\mu}{m}+6\ln\frac{\mu}{m}-L_{y}^{2}+4-\frac{\pi^{2}}{6}\bigg{]}$
$\qquad\qquad-4(1+L_{y})\,\mathcal{L}_{0}(1-z)-8\mathcal{L}_{1}(1-z)+\tilde{\mathcal{R}}(z,b_{T}m)\bigg{\}}\,,$ (3.43)
where we have used the shorthands
$L_{b}$ $\equiv 2\ln\frac{b_{T}\mu}{2e^{-\gamma_{E}}}\,,\qquad y\equiv b_{T}m\,,\quad L_{y}\equiv 2\ln\frac{y}{2e^{-\gamma_{E}}}\,,$ (3.44)
and the regular term ~$\tilde{\mathcal{R}}(y)$ is again integrable for $z\to 0,1$ and given by
~$\tilde{\mathcal{R}}(z,y)$ $\equiv y\int_{0}^{\infty}\!\mathrm{d}t\,J_{1}(ty)\,\mathcal{R}(z,t^{2})$
$=\frac{4}{\bar{z}}\Bigl{[}1+L_{y}+(1+z^{2})K_{0}\Bigl{(}\frac{y\bar{z}}{z}\Bigr{)}-y\bar{z}K_{1}\Bigl{(}\frac{y\bar{z}}{z}\Bigr{)}+2\ln\bar{z}\Bigr{]}\,.$ (3.45)
While we were not able to find a closed-form expression for the total $z$ integral of $\tilde{\mathcal{R}}$, it can readily be obtained numerically or analytically in asymptotic limits through the Hankel transform of eq. (3.3).
3.5 Nonvalence channels at $\mathcal{O}(\alpha_{s})$
Figure -1077: Diagrams contributing to the gluon-initiated partonic heavy-quark TMD FF and the mass-dependent TMD FF matching coefficients at $\mathcal{O}(\alpha_{s})$. Heavy quark propagators and field insertions are indicated in red. The mirror diagrams for (a) and (b) are understood and are included in expressions given in the text.
For completeness we also compute the partonic TMD FFs (or TMD FF matching coefficients onto collinear distributions) for all remaining partonic channels involving heavy quarks. The relevant diagrams at $\mathcal{O}(\alpha_{s})$ are given in figure -1077, where we labeled the identified final-state “hadron” by the momentum $p$. Note that for diagrams (b) and (c) compared to the otherwise identical diagrams in figure -1078, this also changes the definition of $k_{\perp}$, as the coordinate system needs to be adjusted such that the gluon now has vanishing transverse momentum. Results for antiquarks are identical at this order. For the $g\to Q$ contribution, diagram (a) evaluates to
$d_{1\,Q/g}^{(1)}(z,k_{T})=\frac{\alpha_{s}}{4\pi}T_{F}\frac{2}{\pi z^{2}}\frac{k_{T}^{2}z^{4}\bigl{(}2z^{2}-2z+1\bigr{)}+m^{2}z^{2}}{\bigl{(}k_{T}^{2}z^{2}+m^{2}\bigr{)}^{2}}\,.$ (3.46)
For gluon production off the heavy quark, we find after summing over diagrams (b) and (c) and including the mirror diagram for (b),
$\mathcal{J}_{g/Q}^{(1)}(z,k_{T},m)=\frac{\alpha_{s}}{4\pi}C_{F}\frac{2}{\pi}\frac{k_{T}^{2}\bigl{(}z^{2}-2z+2\bigr{)}+m^{2}z^{2}}{z\bigl{(}k_{T}^{2}+m^{2}\bigr{)}^{2}}\,.$ (3.47)
It is straightforward to verify from eq. (3.13) that at $10<z<$ 1 and $k_{T}>0$ the above expression satisfies $1/z^{2}\mathcal{J}_{g/Q}^{(1)}(z,k_{T},m)=d_{1Q/Q}^{(1)}(1-z,k_{T}\frac{z}{1-z})$, as expected from $z\leftrightarrow 1-z$ and the simultaneous change of coordinate system; similarly, $d_{1\,Q/g}^{(1)}(z,k_{T})=d_{1Q/g}^{(1)}(1-z,k_{T}\frac{z}{1-z})$. Note that the bare results in eqs. (3.46) and (3.47) are finite at $z>0$ in $d=4$ and without a rapidity regulator on their own, as expected because there is no cross term with the soft function or the UV renormalization at this order. This is in particular the case as $k_{T}\to 0$, where all singularities in $k_{T}$ are now cut off by the mass for any value of $z>0$.
The effect of the heavy quark on the gluon-onto-gluon matching coefficient at one loop can be written as
$\mathcal{J}_{g/g}^{(1)}(z,k_{T},m,\mu,\zeta)\equiv\mathcal{J}_{g/g}^{(1,n_{\ell})}(z,k_{T},\mu,\zeta)+\Delta\mathcal{J}_{g/g}^{(1,h)}(z,k_{T},m)\,,$ (3.48)
and arises purely from its contribution to the gluon wave function renormalization; here $\mathcal{J}_{g/g}^{(1,n_{\ell})}(z,k_{T})$ is the well-known NLO matching coefficient for the gluon TMD onto the collinear gluon FF in the case where both are defined in a purely light theory with $n_{\ell}$ flavors, see eq. (B). After $\overline{\text{MS}}$-renormalizing the matching coefficient, the net result is
$\Delta\mathcal{J}_{g/g}^{(1)}(z,k_{T},m,\mu)=\frac{\alpha_{s}T_{F}}{4\pi}\,\delta(1-z)\frac{1}{\pi}\delta(k_{T}^{2})\,\frac{4}{3}\ln\frac{m^{2}}{\mu^{2}}\,.$ (3.49)
This agrees with ref. [56], where the same contribution was recently revisited as part of computing all secondary heavy-quark effects in the two-loop gluon TMD PDF; we refer to that reference for a detailed discussion of how it ensures the proper TMD renormalization properties when connecting the theories with $n_{\ell}$ and $n_{\ell}+1$ quark degrees of freedom.
Taking cumulant integrals, we find
$d_{1\,Q/g}^{(1)}(z,{k_{T}^{\mathrm{cut}}})$ $=\frac{\alpha_{s}}{4\pi}T_{F}\frac{2}{z^{2}}\biggl{\{}(2z^{2}-2z+1)\ln(x_{\mathrm{cut}}z^{2}+1)+\frac{2x_{\mathrm{cut}}(1-z)z^{3}}{{x_{\mathrm{cut}}}z^{2}+1}\biggr{\}}\,,$
$\mathcal{J}_{g/Q}^{(1)}(z,{k_{T}^{\mathrm{cut}}},m)$ $=\frac{\alpha_{s}}{4\pi}C_{F}\frac{2}{z}\biggl{\{}(z^{2}-2z+2)\ln(x_{\mathrm{cut}}+1)+\frac{2x_{\mathrm{cut}}(z-1)}{1+x_{\mathrm{cut}}}\biggr{\}}\,,$ (3.50)
where $x_{\mathrm{cut}}\equiv({k_{T}^{\mathrm{cut}}}/m)^{2}$. Finally, taking a $J_{1}$ Hankel transform of the cumulant results as in eq. (3.42), we arrive at the following $b_{T}$-space results:
$d_{1\,Q/g}^{(1)}(z,b_{T})=$ $\frac{\alpha_{s}}{4\pi}T_{F}\frac{4}{z^{2}}\biggl{\{}(2z^{2}-2z+1)K_{0}\Bigl{(}\frac{b_{T}m}{z}\Bigr{)}+b_{T}m(1-z)K_{1}\Bigl{(}\frac{b_{T}m}{z}\Bigr{)}\biggr{\}}\,,$
$\mathcal{J}_{g/Q}^{(1)}(z,b_{T},m)$ $=\frac{\alpha_{s}}{4\pi}C_{F}\frac{4}{z}\bigl{\{}(z^{2}-2z+2)K_{0}(b_{T}m)-b_{T}m(1-z)K_{1}(b_{T}m)\bigr{\}}\,.$ (3.51)
4 bHQET TMD fragmentation factors at NLO
4.1 All-order renormalization properties
In ref. [1] we showed that up to a normalization factor $N_{H}$ involving Clebsch-Gordan coefficients, the unpolarized bHQET TMD fragmentation factor $\chi_{1,H}$ in eq. (2.18) for a heavy hadron state $H$ is given by the trace of the spin density matrix $\rho_{\ell}$ encoding the dynamics of the light constituents $\ell$ of the hadron,
$\chi_{1,H}(b_{T},\mu,\rho)=N_{H}\sum_{h_{\ell}}\rho_{\ell,h_{\ell}h_{\ell}}(b_{\perp},\mu,\rho)\,.$ (4.1)
Here we use the shorthand $\rho\equiv\bar{n}\cdot v$ as introduced below eq. (2.19), which can be interpreted as the total boost of the fragmenting heavy-quark system with velocity $v$ in the frame where the Wilson line direction $\bar{n}$ has unit spatial component. A priori, the matrix element is invariant under boosts – or reparametrization invariance type-III transformations (RPI-III) in SCET [57, 58] – that would take $\bar{n}\to\alpha\,\bar{n}$ for an arbitrary factor $\alpha$, but similarly to the case of relativistic TMD FFs, the presence of rapidity divergences breaks this invariance such that $\chi_{1,H}$ and $\rho_{\ell}$ develop an anomalous dependence on $\rho$. In this subsection, we make the discussion of this point in ref. [1] fully explicit, allowing us to cross check our predictions for the renormalization properties at one loop in section 4.3.
Spelling out all regulators and its open light helicity indices $h_{\ell},h_{\ell}^{\prime}$, the renormalized light-spin density matrix is defined as
$\rho_{\ell,h_{\ell}h_{\ell}^{\prime}}(b_{\perp},\mu,\rho)=\lim_{\epsilon\to 0}Z_{\chi_{1}}^{-1}(\mu,\rho,\epsilon)\lim_{\eta\to 0}\Bigl{[}\rho_{\ell,h_{\ell}h_{\ell}^{\prime}}^{\mathrm{bare}}\Bigl{(}b_{\perp},\epsilon,\eta,\frac{\rho}{\nu}\Bigr{)}\,\sqrt{S^{(n_{\ell})}}(b_{T},\epsilon,\eta,\nu)\Bigr{]}\,,$ (4.2)
where $S^{(n_{\ell})}$ is the bare TMD soft function in the theory where only the $n_{\ell}$ light quark degrees of freedom are active. This is the same soft function as appears in light-quark TMD PDFs or FFs at transverse momenta below the heavy-quark threshold: The fact that it also arises here and cancels the rapidity divergences as $\eta\to 0$ in the bare light-spin density matrix follows from consistency of the TMD factorization for flavor-changing hard processes, e.g. charged-current DIS $e^{-}d\to\nu_{e}\,c$111 We assume as in section 2.1 that a rapidity regulator $\eta$ is used for which soft (zero-bin) subtractions in the bare collinear matrix element are scaleless. If this is not the case, it is crucial that zero-bin subtractions are defined by modes that are soft in a frame where the heavy quark is highly boosted, i.e., their interactions with the heavy-quark sector are encoded in lightlike Wilson lines and the multiplicative zero-bin subtraction factor is obtained by taking $\bar{n}\cdot v\to\infty$. The UV $\overline{\text{MS}}$ renormalization factor $Z_{\chi_{1}}$ likewise follows from consistency and is given by
$Z_{\chi_{1}}^{-1}\bigl{(}\mu,\sqrt{\zeta}/m,\epsilon\bigr{)}$ $=Z_{\mathrm{UV}}(\mu,\zeta,\epsilon)\,Z_{C_{m}}(m,\mu,\zeta,\epsilon)$
$=1-\frac{\alpha_{s}C_{F}}{4\pi}\frac{1}{\epsilon}\Bigl{(}-4\ln\frac{\sqrt{\zeta}}{m}+2\Bigr{)}+\mathcal{O}(\alpha_{s}^{2})\,,$ (4.3)
where $Z_{\mathrm{UV}}$ is the standard UV TMD renormalization factor in eqs. (2.14) and (3.32), and $Z_{C_{m}}$ is the $\overline{\text{MS}}$ renormalization factor of the Wilson coefficient in eq. (2.18) generated by integrating out the heavy quark [43],
$C_{m}^{\mathrm{bare}}(m,\zeta,\epsilon)$ $=Z_{C_{m}}(m,\mu,\zeta,\epsilon)\,C_{m}(m,\mu,\zeta)\,.$ (4.4)
Both the soft function and the UV renormalization are independent of the hadronic states, and thus universal for all possible traces of $\rho_{\ell}$, i.e., they are the same for all possible TMD fragmentation factors $\chi_{1,H}$ (including polarized ones). In particular, they are the same also for the total TMD fragmentation factor $\chi_{1}$ that we consider in section 4.2. Note that while $Z_{\mathrm{UV}}$ and $Z_{C_{m}}$ individually depend on the dimensionful Collins-Soper scale $\zeta=(\bar{n}\cdot P_{H}/z)^{2}$, consistency requires that only the dimensionless combination $\rho=\sqrt{\zeta}/m$ appears in $Z_{\chi_{1}}$ on the left-hand side of eq. (4.1). The fact that $\rho$ is dimensionless also implies that $Z_{\chi_{1}}(\mu,\rho)$ can only subtract single poles in $\epsilon$, but this is indeed precisely what one expects for the renormalization of the product of Wilson lines with a timelike-lightlike cusp [59].
Finally, the bare light-spin density matrix itself is defined as
[…]
×$\qquad\qquad\quad\times\big{\langle}s_{\ell},h_{\ell}^{\prime},f_{\ell};X\big{\rvert}Y^{\dagger}_{v}(0)\,W_{\eta}(0)\big{\rvert}0\big{\rangle}\,,$ (4.14)
where $\lvert s_{\ell},h_{\ell},f_{\ell}\rangle$ indicates a light constituent state $\ell$ of total spin $s_{\ell}$, spin component $h_{\ell}$ along the direction $z^{\mu}\equiv v^{\mu}-\frac{{\bar{n}}^{\mu}}{{\bar{n}}\cdot v}$, and valence flavor content $f_{\ell}$. The Wilson lines $Y_{v}$ along the timelike direction encode the static interactions with the heavy quark,
$Y_{v}(x)\equiv P\Bigl{[}\exp\Bigl{(}\mathrm{i}g\int_{0}^{\infty}\!\mathrm{d}s\,v\cdot A(x+vs)\Bigr{)}\Bigr{]}\,.$ (4.15)
The rapidity regulator only acts on the Wilson lines $W_{\eta}$ along the lightlike direction $\bar{n}$, which are still defined as in eq. (2.13). Here the rapidity regulator is the only source of boost invariance (RPI-III) breaking, which implies that the bare spin density matrix can only depend on the combination $\rho/\nu=\bar{n}\cdot v/\nu$. A priori, it is not obvious that $\rho=\sqrt{\zeta}/m$, as it appears in the renormalization factor in eq. (4.1), should coincide with the $\rho=\bar{n}\cdot v$ that appears in the bare matrix element, but this is in fact guaranteed by the crucial relation
$\sqrt{\zeta}=\bar{n}\cdot P_{H}/z=m\,\bar{n}\cdot v$ (4.16)
that arises from consistently expanding to leading power in $1/m$ [1].
We conclude this section by summarizing the renormalization group equations for the renormalized $\rho_{\ell}$ in $d=4$ dimensions that follow from the above renormalization procedure:
$\mu\frac{\mathrm{d}}{\mathrm{d}\mu}\rho_{\ell,h_{\ell}h_{\ell}^{\prime}}(b_{\perp},\mu,\rho)$ $=\gamma_{\chi_{1}}\bigl{[}\alpha_{s}(\mu),\rho\bigr{]}\,\rho_{\ell,h_{\ell}h_{\ell}^{\prime}}(b_{\perp},\mu,\rho)\,,$
$\rho\frac{\mathrm{d}}{\mathrm{d}\rho}\rho_{\ell,h_{\ell}h_{\ell}^{\prime}}(b_{\perp},\mu,\rho)$ $=\gamma_{\zeta}^{(n_{\ell})}(b_{T},\mu)\,\rho_{\ell,h_{\ell}h_{\ell}^{\prime}}(b_{\perp},\mu,\rho)\,,$ (4.17)
where the $\mu$ anomalous dimension is given by the difference of the standard TMD UV anomalous dimension $\gamma_{\mu}^{q}$ and the anomalous dimension of $C_{m}$, evaluating both at $\zeta=\rho^{2}m^{2}$,
$\gamma_{\chi_{1}}\bigl{[}\alpha_{s}(\mu),\rho\bigr{]}=\gamma_{\mu}^{q}(\mu,\rho^{2}m^{2})-\gamma_{C_{m}}(\mu,m^{2},\rho^{2}m^{2})=\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{(}-8\ln\rho+4\Bigr{)}+\mathcal{O}(\alpha_{s}^{2})\,.$ (4.18)
We stress that when extracting the anomalous dimension at higher orders, care must be taken to perturbatively reexpress $\alpha_{s}^{(n_{\ell}+1)}$ in terms of $\alpha_{s}(\mu)\equiv\alpha_{s}^{(n_{\ell})}(\mu)$.
From the second line of eq. (4.1), we see that $\rho_{\ell}$ develops an anomalous dependence on the boost parameter $\rho$ governed by the Collins-Soper CS kernel $\gamma_{\zeta}^{(n_{\ell})}$ of the theory where only $n_{\ell}$ light quarks are active. For perturbative $b_{T}$, it is given in our normalization by
$\gamma_{\zeta}^{(n_{\ell})}(b_{T},\mu)=\frac{\alpha_{s}C_{F}}{4\pi}\,\biggl{(}-8\ln\frac{b_{T}\mu}{2e^{-\gamma_{E}}}\biggr{)}+\mathcal{O}(\alpha_{s}^{2})+\mathcal{O}(\Lambda_{\mathrm{QCD}}^{2}b_{T}^{2})\,.$ (4.19)
A natural boundary condition for $\rho_{\ell}$, where it is free of large rapidity logarithms, is given by $\rho\sim 1$, i.e., the rest frame of the heavy quark. By contrast, in physical cross sections, the boost is fixed to $\rho=Q/m\gg 1$, i.e., the lab frame in the $e^{+}e^{-}$ case, or the Breit frame for SIDIS. The $\rho$-evolution between those scales resums the associated large rapidity logarithms. As in the relativistic case, the two-dimensional evolution equations satisfy an integrability condition,
$\rho\frac{\mathrm{d}}{\rho}\gamma_{\chi_{1}}\bigl{[}\alpha_{s}(\mu),\rho\bigr{]}=\mu\frac{\mathrm{d}}{\mu}\gamma_{\zeta}^{(n_{\ell})}(\mu,b_{T})=-2\Gamma_{\mathrm{cusp}}\bigl{[}\alpha_{s}(\mu)\bigr{]}$ (4.20)
[…]
4.2 Sum over hadronic states
An important property of the unpolarized bHQET fragmentation factors is that their description greatly simplifies when summing over all hadrons $H$ containing the heavy quark,
$\chi_{1}(b_{T},\mu,\rho)=\sum_{H}\chi_{1,H}(b_{T},\mu,\rho)=\sum_{\ell}\sum_{h_{\ell}}\rho_{\ell}(b_{\perp},\mu,\rho)\,.$ (4.21)
As indicated, this sum in turn is equal to the sum of traces of the light-spin density matrices for all the possible light states $\ell$. Unlike the case of inclusive fragmentation, $\chi_{1}$ does not simply evaluate to unity by probability conservation. However, as in the inclusive case, the complete sum over states can be performed also here, leaving behind a simple vacuum Wilson line correlator [1],
$\chi_{1}^{\mathrm{bare}}(b_{T},\epsilon,\eta,\bar{n}\cdot v)\equiv\frac{1}{N_{c}}\operatorname{Tr}\big{\langle}0\big{\lvert}W_{\eta}^{\dagger}(b_{\perp})\,Y_{v}(b_{\perp})\,Y^{\dagger}_{v}(0)\,W_{\eta}(0)\big{\rvert}0\big{\rangle}\,.$ (4.22)
Its renormalization proceeds exactly as for the whole $\rho_{\ell}$ in eq. (4.2),
$\chi_{1}(b_{T},\mu,\rho)=\lim_{\epsilon\to 0}Z_{\chi_{1}}^{-1}(\mu,\rho,\epsilon)\lim_{\eta\to 0}\Bigl{[}\chi_{1}^{\mathrm{bare}}\Bigl{(}b_{T},\epsilon,\eta,\frac{\rho}{\nu}\Bigr{)}\,\sqrt{S^{(n_{\ell})}}(b_{T},\epsilon,\eta,\nu)\Bigr{]}\,.$ (4.23)
While eqs. (4.22) and (4.23) are valid nonperturbatively and likely exhibit interesting properties at long distances $b_{T}\sim 1/\Lambda_{\mathrm{QCD}}$, an immediate application of their simple operator structure is to evaluate the correlator perturbatively at $b_{T}\ll 1/\Lambda_{\mathrm{QCD}}$, which we do in the next section. As a byproduct of this, we may then read off the universal matching coefficient $C_{1}$ in eq. (2.19) by summing eq. (2.19) over all hadrons $H$ such that $\sum\chi_{H}=1$ drops out,
$\chi_{1}(b_{T},\mu,\rho)=C_{1}(b_{T},\mu,\rho)+\mathcal{O}(\Lambda_{\mathrm{QCD}}^{2}b_{T}^{2})\,.$ (4.24)
Of course, one can also determine $C_{1}$ by a standard matching calculation, e.g. using non-decoupled heavy-quark fields $h_{v}$ in the operators and heavy-quark external states in both the theory at the scale $\mu\sim 1/b_{T}$ and at the low scale $\mu\sim\Lambda_{\mathrm{QCD}}$ to read off the matching coefficient, and the procedure (in pure dimensional regularization) involves equivalent diagrams and leads to the same result.
4.3 One-loop calculation
Figure -1076: Left: Diagrams contributing to the total bHQET TMD fragmentation factor in the perturbative domain. Wilson lines along the lightlike ${\bar{n}}$ direction are indicated by the slanted double lines. The horizontal double lines indicate the Wilson lines along the timelike $v$ direction. The dashed line indicates on-shell cuts after introducing a complete set of states. The mirror diagram for (a) is understood and included in expressions given in the text. Right: Phase-space integrands $\mathcal{A}$ for the two real-emissions diagrams as a function of the gluon emission rapidity $y$.
To evaluate $\chi_{1}$ perturbatively, it is convenient to reinstate a complete sum over states and transition to momentum space to evaluate the bare matrix element,
$\chi_{1}^{\mathrm{bare}}\Bigl{(}k_{T},\epsilon,\eta,\frac{\bar{n}\cdot v}{\nu}\Bigr{)}=\frac{1}{N_{c}}\operatorname{Tr}\int\!\frac{\mathrm{d}\Omega_{2-2\epsilon}}{2\pi}\,k_{T}^{-2\epsilon}\big{\langle}0\big{\lvert}W_{\eta}^{\dagger}\,Y_{v}\big{\rvert}X\big{\rangle}\big{\langle}X\big{\lvert}\bigl{[}\delta^{(2-2\epsilon)}\bigl{(}k_{\perp}^{\mu}+\mathrm{i}\partial_{\perp}^{\mu}\bigr{)}\,Y^{\dagger}_{v}\,W_{\eta}\bigr{]}\big{\rvert}0\big{\rangle}\,,$ (4.25)
in analogy to eq. (3.10). Virtual diagrams with $X$ the vacuum state are scaleless. The two contributing real emission diagrams at $\mathcal{O}(\alpha_{s})$ in Feynman gauge are shown in figure -1076b. The one-gluon Feynman rule for the $Y_{v}$ Wilson line can e.g. be read off from ref. [42], which uses a covariant derivative convention consistent with the one-gluon Feynman rule we use for $W_{\eta}$. Including the mirror diagram for diagram (a), the two diagrams are both given by simple integrals over the rapidity $y$ of the emitted gluon in the heavy-quark rest frame,
$\chi_{1}^{(a)}$ $=\frac{\alpha_{s}C_{F}}{4\pi}\frac{e^{\epsilon\gamma_{E}}}{\Gamma(1-\epsilon)}\frac{\mu^{2\epsilon}}{\pi k_{T}^{2+2\epsilon+\eta}}\Bigl{(}\frac{\bar{n}\cdot v}{\nu}\Bigr{)}^{-\eta}\int\!\mathrm{d}y\,\mathcal{A}^{(a)}(y,\eta)\,,$
$\chi_{1}^{(b)}$ $=\frac{\alpha_{s}C_{F}}{4\pi}\frac{e^{\epsilon\gamma_{E}}}{\Gamma(1-\epsilon)}\frac{\mu^{2\epsilon}}{\pi k_{T}^{2+2\epsilon}}\int\!\mathrm{d}y\,\mathcal{A}^{(b)}(y)\,,$ (4.26)
where the relevant integrands are given by
$\mathcal{A}^{(a)}(y,\eta)=\frac{2e^{-\eta y}}{e^{y}\,\cosh y}\,,\qquad\mathcal{A}^{(b)}(y)=-\frac{1}{\raisebox{-1.00006pt}{$\cosh^{2}y$}}\,,$ (4.27)
and their integrals evaluate to
$\int\!\mathrm{d}y\,\mathcal{A}^{(a)}(y,\eta)=-\frac{2\pi}{\sin(\eta\pi/2)}=-\frac{4}{\eta}+\mathcal{O}(\eta)\,,\qquad\int\!\mathrm{d}y\,\mathcal{A}^{(b)}(y)=-2\,.$ (4.28)
The origin of the rapidity divergence, like the structure of the integrands as a whole, has a rather intuitive interpretation, which we illustrate in the right panel of figure -1076b: Emissions at large positive gluon rapidity $y$ in the rest frame of the heavy quark are strongly suppressed for either contributing diagram, which is the well-known dead-cone effect [28]. For diagram (b), where the static heavy quark $Y_{v}$ both emits and absorbs the gluon, the contribution to the emission probability (dot-dashed blue) peaks at $y=0$ and is suppressed also at large negative rapidities because the contribution must be even in the rest frame. In contrast to this, the contribution from diagram (a), where the emission couples to the $W_{\eta}$ that arises from integrating out the interactions with the anticollinear and soft sectors and is required for gauge invariance, tends to a large constant as $y\to-\infty$ (dashed red), leading to a rapidity divergence. Of course, this precisely corresponds to the central soft momentum region, and thus after introducing the regulator $e^{-\eta y}$ with $\eta<0$ to dampen the behavior as $y\to\infty$ (solid red), the associated pole can cancel against the soft function as usual.
Adding the one-loop soft function in eq. (3.15), which at this order is independent of the matter content of the theory, we find that the $\eta$ poles cancel for arbitrary $\epsilon$ as they must. Taking $\eta\to 0$, we find, using the tree-level result $\chi_{1}^{(0)}(k_{T})=\tfrac{1}{\pi}\delta(k_{T}^{2})$,
$\lim_{\eta\to 0}\Bigl{[}\chi_{1}^{\mathrm{bare}\,(1)}(k_{T},\epsilon,\eta,\frac{\bar{n}\cdot v}{\nu}\Bigr{)}+\tfrac{1}{2}S^{(n_{\ell})\,(1)}(k_{T},\epsilon,\eta,\nu)\Bigr{]}$
$=\frac{\alpha_{s}C_{F}}{4\pi}\frac{e^{\epsilon\gamma_{E}}}{\Gamma(1-\epsilon)}\frac{\mu^{2\epsilon}}{\pi k_{T}^{2+2\epsilon}}\Bigl{(}4\ln\bar{n}\cdot v-2\Bigr{)}$
$=\frac{\alpha_{s}C_{F}}{4\pi}\,\frac{1}{\pi}\Bigl{[}-\frac{\delta(k_{T}^{2})}{\epsilon}+\mathcal{L}_{0}(k_{T}^{2},\mu)+\mathcal{O}(\epsilon)\Bigr{]}\Bigl{(}4\ln\bar{n}\cdot v-2\Bigr{)}\,.$ (4.29)
The remaining pole in $\epsilon$ is precisely canceled by the predicted counter term in eq. (4.1), confirming the renormalization structure we discussed in section 4.1 at one loop. Taking a Fourier transform of the finite remainder, our final result for the renormalized total TMD fragmentation factor in the perturbative domain, and thus for the matching coefficient $C_{1}$, reads
$\chi_{1}(b_{T},\mu,\rho)$ $=C_{1}(b_{T},\mu,\rho)+\mathcal{O}(\Lambda_{\mathrm{QCD}}^{2}b_{T}^{2})$
$=1+\frac{\alpha_{s}C_{F}}{4\pi}(-L_{b})\bigl{(}4\ln\rho-2\bigr{)}+\mathcal{O}(\alpha_{s}^{2})+\mathcal{O}(\Lambda_{\mathrm{QCD}}^{2}b_{T}^{2})\,,$ (4.30)
where $L_{b}$ was defined in eq. (3.44).
5 Consistency checks in the large and small-mass limits
5.1 Large-mass limit
To check the behavior of our final position-space results for $d_{1\,Q/Q}$ in eq. (3.4), we first note that the result is exponentially suppressed at $z<1$ for $y\equiv mb_{T}\gg 1$ as a consequence of eq. (2.17) and the large-argument behavior of the Bessel $K$ functions, i.e., it is given by $\delta(1-z)$ up to a proportionality factor. To verify the prediction from the heavy-quark limit in eq. (2.20), it is then sufficient to compare the total $z$ integral (or any other $z$ moment) of both sides of eq. (2.20). The total $z$ integral yields
$\int\!\mathrm{d}z\,d_{1\,Q/Q}^{(1)}(z,b_{T},\mu,\zeta)$ $=\frac{\alpha_{s}C_{F}}{4\pi}\biggl{\{}-2L_{b}\ln\frac{\zeta}{m^{2}}+4\ln^{2}\frac{\mu}{m}+6\ln\frac{\mu}{m}+4-L_{y}^{2}-\frac{\pi^{2}}{6}$
$\qquad\qquad-\frac{4\pi}{y}+4\mathcal{L}_{0}(y^{2},b_{0}^{2})+12\delta(y^{2})+\mathcal{O}\Bigl{(}\frac{1}{y^{3}}\Bigr{)}\biggr{\}}\,.$ (5.1)
Here we used eq. (3.3) for the total $z$ integral of the cumulant-space remainder term $\mathcal{R}(z,{k_{T}^{\mathrm{cut}}}/m)$, which we then power expanded for small ${k_{T}^{\mathrm{cut}}}/m$, letting us evaluate the total $z$ integral of eq. (3.4) power by power in $1/m$. We find that the leading $\mathcal{O}(1/m^{0})$ term on the first line is in full agreement with the bHQET prediction from eq. (2.20), using our one-loop result for $C_{1}$ in eq. (4.3) and the known result for $C_{m}$ in eq. (B.2). We point out the interesting observation that the first subleading $\mathcal{O}(1/m)$ correction in eq. (5.1) is free of large logarithms $L_{y}\equiv 2\ln\tfrac{y}{b_{0}}$, $b_{0}\equiv 2e^{-\gamma_{E}}$, suggesting a simple structure if factorized in terms of subleading $\mathcal{O}(1/m)$ TMD matrix elements in bHQET. On the other hand, the terms at the next order require plus regularization in $b_{T}$ space due to their $1/y^{2}$ scaling and do feature a single-logarithmic term $\mathcal{L}_{0}(y^{2},b_{0}^{2})$.
For completeness, we may also expand the nonvalence results in eq. (3.5) for $y\gg 1$ at finite values of $10<z<$ 1 in a similar way, which simply yields
$d_{1\,Q/g}^{(1)}(z,b_{T})$ $=\frac{\alpha_{s}T_{F}}{4\pi}\frac{1}{z^{2}}\biggl{\{}8z^{2}\,\delta(y^{2})+\mathcal{O}\Bigl{(}\frac{1}{y^{3}}\Bigr{)}\biggr{\}}\,,$
$\mathcal{J}_{g/Q}^{(1)}(z,b_{T},m)$ $=\frac{\alpha_{s}C_{F}}{4\pi}\biggl{\{}8z\,\delta(y^{2})+\mathcal{O}\Bigl{(}\frac{1}{y^{3}}\Bigr{)}\biggr{\}}\,,$ (5.2)
i.e., both channels are in fact suppressed by $1/m^{2}$ at finite $z$.
5.2 Small-mass limit
Expanding eqs. (3.4) and (3.5) for small values of the mass, $y\equiv mb_{T}\ll 1$, we obtain:
$d^{(1)}_{1\,Q/Q}(z_{H},b_{T},\mu,\zeta)$ $=\frac{\alpha_{s}C_{F}}{4\pi}\frac{1}{z^{2}}\bigg{\{}\delta(1-z)\bigg{[}-2L_{b}\ln\frac{\zeta}{m^{2}}+4\ln^{2}\frac{\mu}{m}+6\ln\frac{\mu}{m}-L_{y}^{2}+4-\frac{\pi^{2}}{6}\bigg{]}$
$\qquad\qquad\qquad-4(1+L_{y})\,\mathcal{L}_{0}(1-z)-8\mathcal{L}_{1}(1-z)$
$\qquad\qquad\qquad+4+2L_{y}(1+z)+\frac{8\ln\bar{z}-4(1+z^{2})\ln\tfrac{\bar{z}}{z}}{\bar{z}}$
$\qquad\qquad\qquad+y^{2}\biggl{(}\frac{1}{z^{2}}-z-\frac{L_{y}(1+z)^{2}\bar{z}}{2z^{2}}-\frac{(1+z)^{2}\bar{z}\ln\tfrac{\bar{z}}{z}}{z^{2}}\biggr{)}+\mathcal{O}(y^{4})\bigg{\}}\,,$
$d^{(1)}_{1\,Q/g}(z_{H},b_{T},\mu,\zeta)$ $=\frac{\alpha_{s}T_{F}}{4\pi}\biggl{\{}-2L_{y}[2(z-1)z+1]-4(z-1)z+4[2(z-1)z+1]\ln z$
$\qquad\qquad\qquad+y^{2}\biggl{(}-\frac{L_{y}(1-2z)^{2}}{2z^{2}}+\frac{3(z-1)z+1}{z^{2}}+\frac{(1-2z)^{2}\ln z}{z^{2}}\biggr{)}+\mathcal{O}(y^{4})\biggr{\}}\,,$
$\mathcal{J}^{(1)}_{g/Q}(z_{H},b_{T},\mu,\zeta)$ $=\frac{\alpha_{s}C_{F}}{4\pi}\biggl{\{}-\frac{2L_{y}[(z-2)z+2]}{z}+\frac{4(z-1)}{z}$
$\qquad\qquad\qquad+y^{2}\left(-\frac{L_{y}(z-2)^{2}}{2z}+z+\frac{3}{z}-3\right)+\mathcal{O}(y^{4})\biggr{\}}\,.$ (5.3)
It is straightforward to verify that the leading $\mathcal{O}(y^{0})$ terms are in full agreement with the prediction in eq. (2.4) using the one-loop ingredients collected in our notation in Appendix B. At leading power, the logarithms $L_{y}\equiv 2\ln\tfrac{y}{2e^{-\gamma_{E}}}$ in all channels are predicted by the timelike DGLAP evolution between the TMD matching coefficients at $\mu\sim 1/b_{T}$ and the matching kernels and partonic collinear FFs at $\mu\sim m$. The $\mathcal{O}(y^{2})$ terms, which we here give explicitly, likewise feature single logarithms $L_{y}$ in all partonic channels, and can be used to investigate the $\mathcal{O}(b_{T}^{2})$ twist-4 matching of TMD FFs in a physics setup where the low scale matrix elements at $\mu\sim m$ are perturbatively calculable.
6 Joint threshold and $q_{T}$ resummation for massive quark fragmentation
Figure -1075: Relevant modes contributing to the heavy-quark TMD FF as the collinear momentum fraction $z_{H}\to 1$ while counting $m\sim k_{T}$. See the text for a detailed discussion.
The remarkable simplicity of our final $b_{T}$-space result in eq. (3.4) as $z\to 1$ warrants further investigation. Specifically, the leading terms at $\mathcal{O}[(1-z)^{-1}]$ feature only a logarithmic dependence on the mass and $b_{T}$, while in general, a power-like dependence on $mb_{T}\sim 1$ would be allowed, suggesting that the heavy-quark TMD FF can be factorized further in this regime such that the logarithms become predicted by RGEs. To investigate this question, we modify our previous assumption on the power counting of $\bar{z}_{H}\equiv 1-z_{H}$, eq. (2.17), and instead count
[…]
We stress that we continue to count $m\sim k_{T}\sim 1/b_{T}$. In figure -1075 we illustrate the contributing modes in the EFT for this scenario. Figure -1075 (a) shows the original, unmodified mode setup that led to eq. (2.1). Here the partonic heavy-quark TMD FF arises from matching collinear modes (green), accounting for the usual contribution of the central soft modes (orange), onto the nonperturbative modes (gray) describing the brown muck, which is isotropic in the rest frame of the heavy quark indicated by the dashed black line. The collinear modes are kinematically unconstrained in the longitudinal direction because $\bar{z}_{H}\sim 1$.
Once the modified assumption eq. (6.1) is made, collinear modes at the scale $\mu\sim k_{T}\sim m$ can no longer be radiated into the final state. Their effect becomes purely virtual and is encoded in the matching coefficient $C_{m}$ (green cross) from matching onto bHQET modes at yet lower virtuality. Real emissions into the $n$-collinear direction at the scale $\mu\sim k_{T}\sim m$ can instead only have collinear-soft (pink) scaling
$p_{\mathrm{cs}}\sim\Bigl{(}\bar{z}_{H}Q,\tfrac{k_{T}^{2}}{\bar{z}_{H}Q},k_{T}\Bigr{)}\,,$ (6.2)
which is the characteristic mode of double-differential $z\to 1$ threshold and $q_{T}$ resummation, often simply dubbed joint resummation [60, 61, 62, 63, 64, 65, 66, 67, 68].222 We note that while ref. [64] also dealt with final-state heavy quarks, that work considered the joint resummation for the total transverse momentum of a heavy quark pair with respect to the beam direction, which is a problem that is distinct from the one we consider here.
The physical picture at scales $\mu\ll k_{T}\sim m$ in this scenario, which is governed by one or several bHQET modes at their characteristic lab-frame rapidity $p^{-}/p^{+}\sim Q^{2}/m$, further depends on the relative size of $\Lambda_{\mathrm{QCD}}/m$ and $\bar{z}_{H}$, specifically whether $\Lambda_{\mathrm{QCD}}/m\sim\bar{z}_{H}$ or $\Lambda_{\mathrm{QCD}}/m\ll\bar{z}_{H}$. (Since $\Lambda_{\mathrm{QCD}}$ introduces a fourth scale into the problem, i.e., a third independent dimensionless power counting parameter, either case may hold in general and independent of $k_{T}/Q$ and $m/Q$.) The relevant modes for the case $\Lambda_{\mathrm{QCD}}/m\sim\bar{z}_{H}$ are illustrated in figure -1075 (c), and are in fact closely analogous to the well-understood case of single-inclusive heavy-hadron production near the endpoint in $e^{+}e^{-}$ collisions. Specifically, the nonperturbative fragmentation process at the scale $\mu\sim\Lambda_{\mathrm{QCD}}\sim\bar{z}_{H}m$ provides a leading contribution to the measured longitudinal hadron momentum in this case (but not to its transverse momentum, where the contribution is power suppressed!). This effect is encoded in the well-known collinear bHQET fragmentation shape function $F_{H}(k^{-},\mu)$ [5, 6]. [^3] We use the symbol $F_{H}$ here to distinguish the shape function from the TMD soft function; in ref. [6], the symbol $S_{H/Q}$ is used instead, and our partonic $\hat{F}$ corresponds to their $S_{Q/Q}$. In total, the factorization of the hadron-level heavy-quark TMD FF in this regime reads, with $\zeta=\omega^{2}$ the Collins-Soper scale,
$D_{1\,H/Q}(z_{H},b_{T},\mu,\omega^{2})$ $=C_{m}(m,\mu,\omega^{2})\,\int\!\mathrm{d}\ell^{-}\,\mathcal{S}\Bigl{(}b_{T},m,\mu,\frac{\ell^{-}}{\nu}\Bigr{)}\,\sqrt{S}(b_{T},m,\mu,\nu)$
×$\quad\times\int\!\frac{\mathrm{d}k^{-}}{\bar{n}\cdot v}\,F_{H}\Bigl{(}\frac{k^{-}}{\bar{n}\cdot v},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}_{H}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}$
×$\quad\times\biggl{[}1+\mathcal{O}(\bar{z}_{H})+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{k_{T}}\Bigr{)}+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{m}\Bigr{)}\biggr{]}\,,$ (6.3)
where $\mathcal{S}$ and $\sqrt{S}$ are the separately renormalized collinear-soft function from joint resummation [68] and the square root of the standard TMD soft function, respectively. By evolving all terms in eq. (6) from their canonical scales (e.g. in Mellin space $N\leftrightarrow z_{H}$ where $\nu_{\mathcal{S}}\sim\bar{n}\cdot v\,\mu_{F_{H}}\sim Q/N$ and the evolution is multiplicative) to the common scale $\mu\sim\sqrt{\zeta}\sim Q$ with the hard function, all large threshold and transverse momentum logarithms present in the heavy-hadron TMD cross section in this regime can be resummed simultaneously.
It is instructive to compare the above to the refactorization of the light-parton TMD FF in the joint resummation regime $k_{T}\ll\bar{z}_{h}Q\ll Q$, which follows from crossing the light-parton TMD PDF refactorization relations [68],
$D_{1\,h/q}(z_{h},b_{T},\mu,\omega^{2})$ $=\int\!\mathrm{d}\ell^{-}\,\mathcal{S}\Bigl{(}b_{T},m,\mu,\frac{\ell^{-}}{\nu}\Bigr{)}\,\sqrt{S}(b_{T},m,\mu,\nu)$
×$\quad\times\int\!\mathrm{d}k^{-}\,D^{\mathrm{thr}}_{h/q}\Bigl{(}\frac{k^{-}}{P_{h}^{-}},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}_{h}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}$
×$\quad\times\biggl{[}1+\mathcal{O}(\bar{z}_{H})+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{k_{T}}\Bigr{)}+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{m}\Bigr{)}\biggr{]}\,,$ (6.4)
where $D^{\mathrm{thr}}_{h/q}$ is the collinear (threshold) fragmentation function, whose renormalization is given by the timelike DGLAP kernels expanded to leading power in $1-z$. The product $C_{m}\,F_{H}$ in eq. (6) is renormalized like a threshold FF [6], thus eq. (6) is indeed RGE consistent by comparison with eq. (6). We note that both eqs. (6) and (6) generically feature a dependence on the heavy quark mass in the soft and collinear-soft function since $mb_{T}\sim 1$, to which we return towards the end of this section.
To make contact with the partonic heavy-quark TMD FF, we next take $\Lambda_{\mathrm{QCD}}\to 0$, which in particular means that now $\Lambda_{\mathrm{QCD}}/m\ll\bar{z}_{H}$. The mode setup for this regime is illustrated in figure -1075 (b). The only change compared to figure -1075 (c) is that the fragmentation shape function now factorizes into the perturbative bHQET shape function $\hat{F}$ and the total fragmentation probability $\chi_{H}$ [6],
$F_{H}(k^{-},\mu)=\hat{F}(k^{-},\mu)\,\chi_{H}+\mathcal{O}(\Lambda_{\mathrm{QCD}})\,.$ (6.5)
This implies that the joint resummation of the partonic heavy-quark TMD FF is governed by the following refactorization relation,
$d_{1\,Q/Q}(z_{H},b_{T},\mu,\omega^{2})$ $=C_{m}(m,\mu,\omega^{2})\,\int\!\mathrm{d}\ell^{-}\,\mathcal{S}\Bigl{(}b_{T},m,\mu,\frac{\ell^{-}}{\nu}\Bigr{)}\,\sqrt{S}(b_{T},m,\mu,\nu)$
×$\quad\times\int\!\frac{\mathrm{d}k^{-}}{\bar{n}\cdot v}\,\hat{F}\Bigl{(}\frac{k^{-}}{\bar{n}\cdot v},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}_{H}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}+\mathcal{O}\bigl{[}(1-z_{H})^{0}\bigr{]}\,.$ (6.6)
This is precisely the relation governing the leading $1/\bar{z}$ terms in the partonic heavy-quark TMD FF that we were seeking out, and it is straightforward to verify eq. (6) at one-loop order against eq. (3.4) using the fact that ~$\tilde{\mathcal{R}}(z,y)=\mathcal{O}[(1-z)^{0}]$, the fact that $\bar{n}\cdot v=\omega/m$, and the ingredients given in appendix B. In particular, since the soft and soft-collinear function are defined as Wilson line correlators, they can only acquire mass dependence through secondary mass effects starting at $\mathcal{O}(\alpha_{s}^{2}C_{F}T_{F})$. In turn, this implies that the mass dependence at $\mathcal{O}(\alpha_{s})$ arises only from that of $C_{m}$, which is logarithmic to all orders, explaining why eq. (3.4) has to have the simple logarithmic mass dependence that we found for $\bar{z}\to 1$. We stress that starting at $\mathcal{O}(\alpha_{s}^{2})$, the soft and soft-collinear functions do develop a highly nontrivial dependence on $mb_{T}$, which is e.g. inherited by the coefficients of $\mathcal{L}_{0}(\bar{z})$ and $\delta(\bar{z})$ in $d_{1\,Q/Q}^{(2)}$ at two loops. The required ingredients are readily available in the literature [44], [^4] The secondary mass effects in $\sqrt{S}$ have been calculated in ref. [44]. For the collinear-soft function one may use that for the exponential regulator of ref. [67], the collinear-soft function is simply given by an exponential of the Collins-Soper kernel to all orders [69]. This remains true in the presence of mass effects in the CS kernel, which have likewise been calculated in ref. [44]. One potential subtlety is that the results of ref. [44] for $\sqrt{S}$ were obtained using the $\eta$ regulator of refs. [39, 40] instead. However, it is easy to see that the one-loop soft function with a massive gluon (which underlies the dispersive calculation of the secondary quark mass effects in ref. [44]) leads to identical results for both regulators, meaning the final renormalized result for the mass-dependent soft function in ref. [44] is also valid for either regulator. which can provide a strong cross check on a future two-loop calculation of the partonic heavy-quark TMD FF, or simplify it by allowing to restrict to finite $z<1$ from the start.
7 Application: EEC with heavy quarks in the back-to-back limit
7.1 Mass-dependent EEC jet functions
In this section we use our results to produce explicit results for the the effect of heavy-quarks on the energy-energy correlator (EEC) observable [70] as measured in $e^{+}e^{-}$ collisions,
$\mathrm{EEC}(\chi)\equiv\frac{1}{\sigma_{\mathrm{total}}}\,\frac{\mathrm{d}\sigma}{d\cos\chi}=\frac{1}{\sigma_{\mathrm{total}}}\sum_{a,b\,\in X}\int\!\mathrm{d}\sigma_{ee\to X}\frac{E_{a}E_{b}}{Q^{2}}\delta\bigl{[}\cos(\theta_{ab})-\cos(\chi)\bigr{]}\,,$ (7.1)
where the sum runs over all pairs of hadrons $a$ and $b$ in the event with energies $E_{a},E_{b}$ and $Q$ is the center-of-mass energy. We are specifically interested in the so-called back-to-back limit $\chi\to\pi$, where a parton-level factorization theorem has been proven [21] and can be directly related to TMD factorization. This precisely complements the study of ref. [71], where the NLO mass effects in the opposite (collinear) limit $\chi\to 0$ of the EEC were calculated, and can serve as data for a future calculation of the heavy-quark EEC for general angles.
To do so, it is convenient to rewrite the EEC as a differential distribution in $z_{\chi}=\bigl{[}1-\cos(\theta_{ab})\bigr{]}/2\to 1$, i.e., the average pairwise angle of energy-weighted hadron pairs, $\mathrm{EEC}(\chi)=(1/\sigma_{\mathrm{total}})\,\mathrm{d}\sigma/\mathrm{d}z_{\chi}$. We are interested in the modifications of the fixed-order EEC from the presence of heavy quarks in this limit, counting $(1-z_{\chi})Q^{2}\sim 1/b_{T}^{2}\sim m^{2}$, where $b_{T}$ – precisely as in TMD factorization – is the typical transverse spacetime separation of fields in the relevant correlators. The $z_{\chi}$-differential distribution is given by $z_{\mathfrak{h}}$-moments of the TMD FFs for identified hadrons [21],
$\frac{1}{\sigma_{0}}\frac{\mathrm{d}\sigma}{\mathrm{d}z_{\chi}}$ $=\sum_{i,j}\mathcal{H}_{ee\to ij}(Q^{2},\mu)~{}Q^{2}\!\!\int_{0}^{\infty}\!\mathrm{d}b_{T}\,b_{T}\,J_{0}\bigl{(}\sqrt{1-z_{\chi}}\,b_{T}Q\bigr{)}$ (7.2)
×$\quad\times\sum_{\mathfrak{h},\mathfrak{h}^{\prime}}\int\!\mathrm{d}z_{\mathfrak{h}}\,z_{\mathfrak{h}}^{3}D_{1\,\mathfrak{h}/i}(z_{\mathfrak{h}},b_{T},\mu,Q^{2})\int\!\mathrm{d}z_{\mathfrak{h}^{\prime}}\,z_{\mathfrak{h}^{\prime}}^{3}D_{1\,\mathfrak{h}^{\prime}/j}(z_{\mathfrak{h}^{\prime}},b_{T},\mu,Q^{2})$
$\equiv\sum_{i,j}\mathcal{H}_{ee\to ij}(Q^{2},\mu)~{}Q^{2}\!\!\int_{0}^{\infty}\!\mathrm{d}b_{T}\,b_{T}\,J_{0}\bigl{(}\sqrt{1-z_{\chi}}\,b_{T}Q\bigr{)}$
×$\quad\times J_{i}(b_{T},m,\mu,Q/\nu)\,J_{j}(b_{T},m,\mu,Q/\nu)\,S(b_{T},m,\mu,\nu)\,,$
where the sum runs over all possible pairs of partons $i,j$, with nonzero contributions for $i,j=q,\bar{q}$ a light quark pair or $i,j=Q,\bar{Q}$ a heavy quark pair, and it is understood that $\mathcal{H}_{ee\to gg}=0$. Here our normalization is such that $\mathrm{d}\sigma/\mathrm{d}z_{\chi}=\sigma_{0}\,\delta(1-z_{\chi})+\mathcal{O}(\alpha_{s})$ in the limit where the process is mediated only by virtual photons, and the hard function $\mathcal{H}_{ee\to ij}$ is given in appendix B. On the second equality we defined the EEC back-to-back jet function $J_{i}(b_{T},m,\mu,\nu)$ for a – possibly massive – parent parton $i$, and we followed the usual convention in the EEC literature to factor off a rapidity and UV-renormalized TMD soft function $S(b_{T},m,\mu,\nu)$. Both $J_{i}$ and $S$ acquire mass dependence for $1/b_{T}^{2}\sim m^{2}$ from primary and secondary heavy-quark effects, which we now classify and evaluate further. Note that while the energy fractions entering eq. (7.1) in general differ from the lightcone momentum fractions $z_{\mathfrak{h}}$ that the TMD FFs depend on, in particular for massive particles, the difference is power-suppressed in the back-to-back limit where $k_{T},m,m_{\mathfrak{h}}\ll z_{\mathfrak{h}}Q$.
To evaluate the sums over hadrons, we break each of them down into two terms, a sum over light hadrons $\mathfrak{h}=h$ and one over heavy hadrons $\mathfrak{h}=H$. For light-hadron final states, the standard manipulations of the collinear FF using its momentum sum rule apply [21], such that also in the presence of intermediate heavy quarks and for all of $i=q,Q,g$,
$\sum_{h}\int\!\mathrm{d}z_{h}\,z_{h}^{3}\,D_{1\,h/i}(z_{h},b_{T},\mu,Q^{2})$
$=\sum_{k=q,\bar{q},g}\int\!\mathrm{d}z\,z\,\mathcal{J}_{k/i}(z,b_{T},m,\mu,Q^{2})+\mathcal{O}(\Lambda_{\mathrm{QCD}}^{2}b_{T}^{2})\,,$ (7.3)
where $\mathcal{J}_{1\,k/i}$ is the mass-dependent matching coefficient introduced in eq. (2.16). Importantly, at the scale $\mu\sim\Lambda_{\mathrm{QCD}}$ light partons can only fragment to light hadrons in the theory with $n_{\ell}$ light quark flavors, which means that the collinear FF momentum sum rule holds for the set of light hadrons $\{h\}$ on their own within that theory.
On the other hand, for heavy-hadron final states $H$ containing either $Q$ or $\bar{Q}$, we have
$\sum_{H}\int\!\mathrm{d}z_{H}\,z_{H}^{3}\,D_{1\,H/i}(z_{H},b_{T},\mu,Q^{2})$
$=\sum_{K=Q,\bar{Q}}\int\!\mathrm{d}z\,z^{3}\,d_{1\,K/i}(z,b_{T},\mu,Q^{2})+\mathcal{O}(\Lambda_{\mathrm{QCD}}^{2}b_{T}^{2})\,,$ (7.4)
which again holds for all of $i=q,Q,g$. While the end result looks similar to eq. (7.1), the moment of the perturbative coefficient arises in a different (and much simpler) way, namely because the fragmentation probabilities $\chi_{H}$ from eq. (2.1) sum to unity, while $d_{1\,Q/i}$ and $d_{1\,\bar{Q}/i}$ are independent of the state $H$. We note in passing that bound states containing multiple heavy quarks also in general contribute to the sum over hadrons in eq. (7.2) and e.g. for a $Q\bar{Q}$ bound state can be evaluated in terms of the TMD FF matching coefficients onto NRQCD long-distance matrix elements calculated in refs. [9, 10, 11], but this contribution to the jet function $J_{i}$ only starts at $\mathcal{O}(\alpha_{s}^{2})$.
The expressions above are particular ($N=1$) Mellin moments of the partonic heavy-quark TMD FF and the TMD FF matching coefficients, which we in general define as
$d_{1\,Q/i}^{[N]}(b_{T},\mu,\zeta)$ $\equiv\int\!\mathrm{d}z\,z^{N+2}d_{1\,Q/i}(z,b_{T},\mu,\zeta)\,,$
$\mathcal{J}_{k/i}^{[N]}(b_{T},m,\mu,\zeta)$ $\equiv\int\!\mathrm{d}z\,z^{N}\mathcal{J}_{k/i}(z,b_{T},m,\mu,\zeta)\,,$ (7.5)
where for $d_{1\,Q/i}$ we explicitly compensate for the usual measure factor of $z^{2}$ when taking moments at fixed $b_{T}$ (or $k_{T}$). In summary, we have
$J_{i}(b_{T},m,\mu,Q/\nu)\bigl{[}S(b_{T},m,\mu,\nu)\bigr{]}^{1/2}$
$=\sum_{K=Q,\bar{Q}}d_{1\,K/i}^{[1]}(b_{T},\mu,Q^{2})+\sum_{k=q,\bar{q},g}\mathcal{J}_{k/i}^{[1]}(b_{T},m,\mu,Q^{2})\,,$ (7.6)
for the quark, massive quark, or gluon EEC back-to-back jet functions, which generalizes the defining relation for the massless jet function in ref. [21]. We collect the Mellin moments of our new results that enter the right-hand side in appendix C. We point out that the organization of our final one-loop result for the heavy-quark TMD FF in momentum space in terms of plus distributions on rectangular domains in $(z,k_{T})$, see eq. (3.2), makes it particularly straightforward to take moments with respect to $z$.
7.2 Jet function consistency in the small and large-mass limits
The all-order consistency relation for the mass-dependent EEC jet functions in the small-mass limit follows from eq. (2.4) and reads
$J_{i}(b_{T},m,\mu,Q/\nu)=J_{i}(b_{T},\mu,Q/\nu)+\mathcal{O}(mb_{T})\,,$ (7.7)
i.e., they simply reduce to their massless counterparts because the partonic heavy-quark collinear FFs and the FF flavor matching kernels preserve the momentum sum rule. (In other words, the measurement becomes fully inclusive over the details of the IR also in this case, which now includes the mass.)
On the other hand, in the large-mass limit we can infer from the discussion in section 2.3 that the heavy-quark jet function to all orders satisfies
$J_{Q}(b_{T},m,\mu,Q/\nu)\bigl{[}S(b_{T},m,\mu,\nu)\bigr{]}^{1/2}=C_{m}(m,\mu,Q^{2})\,C_{1}\Bigl{(}b_{T},\mu,\frac{Q}{m}\Bigr{)}\,.$ (7.8)
Note that it is not power suppressed in this regime because even at scales below the quark mass, the heavy quark continues to radiate light modes and pick up recoil, leading to a nontrivial angular dependence, and thus a nonzero EEC. As for the TMD FFs, the mass effects in the light-parton jet functions $J_{q}$ and $J_{g}$ for $(1-z_{\chi})Q^{2}\ll m^{2}$ are of purely virtual origin and encoded in collinear and soft mass matching coefficients onto their counterparts in the $n_{\ell}$ (purely light) theory, see refs. [44, 56].
7.3 Complete mass dependence of the EEC at $\mathcal{O}(\alpha_{s})$
The mass effects in the light-quark mediated cross section $i=q,\bar{q}$ only enter through closed heavy-quark loops or pair production starting at $\mathcal{O}(\alpha_{s}^{2})$, such that at $\mathcal{O}(\alpha_{s})$ the total mass effect simply reads
$\frac{1}{\sigma_{0}}\frac{\mathrm{d}\sigma^{(1)}}{\mathrm{d}z_{\chi}}$ $=\frac{1}{\sigma_{0}}\frac{\mathrm{d}\sigma^{(1)}_{ee\to q\bar{q}}}{\mathrm{d}z_{\chi}}+2\mathcal{H}_{ee\to Q\bar{Q}}^{(1)}(Q^{2},\mu)\,\delta(1-z_{\chi})$ (7.9)
$\quad+4\mathcal{H}_{ee\to Q\bar{Q}}^{(0)}\,\pi Q^{2}\Bigl{\{}d_{1\,Q/Q}^{[1]\,(1)}\bigl{(}\sqrt{1-z_{\chi}}\,Q,\mu,Q^{2}\bigr{)}+\mathcal{J}_{g/Q}^{[1]\,(1)}\bigl{(}\sqrt{1-z_{\chi}}\,Q,\mu,Q^{2}\bigr{)}\Bigr{\}}\,,$
where $\mathrm{d}\sigma_{ee\to q\bar{q}}$ refers to the light quark-mediated cross section, which is known to $\mathcal{O}(\alpha_{s}^{3})$ [72], and the expression for the hard function $\mathcal{H}_{ee\to Q\bar{Q}}$ is given in appendix B. We also exploited that the convolution (or the inverse Fourier transform) is trivial at fixed $\mathcal{O}(\alpha_{s})$, and that the contributing partonic TMD FFs are all equal. Using the results of appendix C for $N=1$ and adding the contribution from the one-loop hard-function, we find
$\frac{1}{\sigma_{0}}\frac{\mathrm{d}\sigma^{(1)}}{\mathrm{d}z_{\chi}}$ $=\frac{1}{\sigma_{0}}\frac{\mathrm{d}\sigma^{(1)}_{ee\to q\bar{q}}}{\mathrm{d}z_{\chi}}+2\mathcal{H}_{ee\to Q\bar{Q}}^{(0)}\,\frac{\alpha_{s}C_{F}}{4\pi}\biggl{\{}8(2\ln\rho-1)\,\mathcal{L}_{0}(\bar{z}_{\chi})+8\Bigl{(}2\ln^{2}\rho+\ln\rho+\frac{2\pi^{2}}{3}-2\Bigr{)}\,\delta(\bar{z}_{\chi})$
$\qquad+4\rho^{2}\frac{\pi(-3-18x+x^{2})+\sqrt{x}(9+x-9x^{2}-x^{3})-\sqrt{x}(1+24x+9x^{2}+2x^{3})\ln x}{\sqrt{x}(1+x)^{4}}\biggr{\}}$
$\quad+\mathcal{O}(\alpha_{s}^{2})+\mathcal{O}\bigl{(}\bar{z}_{\chi}^{0}\bigr{)}+\mathcal{O}\Bigl{(}\frac{1}{\rho}\Bigr{)}\,,$ (7.10)
where we used the shorthands $\rho\equiv Q/m$ for the boost of the heavy quarks, $\bar{z}_{\chi}\equiv 1-z_{\chi}$, and $x\equiv\rho^{2}\bar{z}_{\chi}=k_{T}^{2}/m^{2}$. The result is valid up to power corrections in $1/Q$, i.e., $\mathcal{O}(\bar{z}_{\chi}^{0})$ or $\mathcal{O}(1/\rho)$, but captures the exact dependence on $\rho^{2}\bar{z}_{\chi}\sim 1$.
7.4 Extension to nonperturbative $1-z_{\chi,H}\sim\Lambda_{\mathrm{QCD}}^{2}/Q^{2}$
To close this section, we note that an intriguing modification to the observable can be made where the sum in the observable definition in eq. (7.1) is restricted to heavy hadrons $a,b\in\{H\}$ containing valence heavy quarks. For this modified observable, which we may call $\mathrm{EEC}_{H}$ with an associated differential variable $z_{\chi,H}$, it is in fact straightforward to extend the factorization all the way down into the nonperturbative region $1-z_{\chi,H}\sim\Lambda_{\mathrm{QCD}}^{2}/Q^{2}$:
$\frac{1}{\sigma_{0}}\frac{\mathrm{d}\sigma}{\mathrm{d}z_{\chi,H}}$ $=\mathcal{H}_{ee\to Q\bar{Q}}(Q^{2},\mu)\,C_{m}^{2}(m,\mu,Q^{2})$
×$\quad\times Q^{2}\!\!\int_{0}^{\infty}\!\mathrm{d}b_{T}\,b_{T}\,J_{0}\bigl{(}\sqrt{1-z_{\chi,H}}\,b_{T}Q\bigr{)}\Bigl{[}\chi_{1}\Bigl{(}b_{T},\mu,\frac{Q}{m}\Bigr{)}\Bigr{]}^{2}\,,$ (7.11)
where $\chi_{1}$ is the renormalized Wilson-line correlator defined in eq. (4.23). The simplicity of eq. (7.4), which is valid for $\Lambda_{\mathrm{QCD}}\lesssim\sqrt{1-z_{\chi,H}}\,Q\ll m\ll Q$ and coincides with eq. (7.8) for $\Lambda_{\mathrm{QCD}}\ll\sqrt{1-z_{\chi,H}}\,Q$, comes about exactly as for the individual $\chi_{1}$ in section 4: Summing over heavy hadrons $H$ allows one to complete the sum over states in the EFT and reduces the final result to (a square of) vacuum correlators of Wilson lines. We note that in ref. [1], we already mentioned the factorization in eq. (7.4) as a factorization theorem for the total heavy-hadron TMD cross section in $e^{+}e^{-}$ collisions, but that – in hindsight – of course coincides with the EEC in the back-to-back limit.
8 Comparison to 2310.19207 by Dai, Kim, and Leibovich
Recently, an independent NLO calculation of the heavy-quark TMD FFs $d_{1\,Q/Q}$ and $d_{1\,Q/g}$ was performed in ref. [13]. Before commenting on the differences between the two calculational approaches, we first want to stress that after accounting for a measure factor of $1/z^{2}$ and the renormalized TMD soft function in eq. (B.4), we find complete agreement of our eqs. (3.4) and (3.5) with the final $b_{T}$-space results given in their eqs. (2.36) and (4.26). In order to gauge the strength of this cross check, it is relevant to note that while the calculation of the real-emission diagrams in ref. [13], like ours in section 3.1, was initially performed in momentum space and only Fourier-transformed after renormalization, ref. [13] used a completely orthogonal method to isolate the singularities in $d^{(1)}_{1\,Q/Q}$, i.e., to address the key source of subtleties. Specifically, in ref. [13] the characteristic denominator $1/X\equiv 1/(k_{T}^{2}+m^{2}\bar{z}^{2}/z^{2})$ appearing in the real-emission diagrams was directly expanded into one-dimensional plus distribution of $X$ and in addition was multiplied by one-dimensional distributions of $1-z$, which effectively amounts to performing a change of variables to $(z,X)$. This procedure is completely orthogonal to our approach, namely, we expand in terms of one-dimensional plus distributions along the boundaries $\bar{z}=0$ and $k_{T}=0$ and a two-dimensional plus distribution whose defining domain aligns with the main variables $(z,k_{T})$. The two approaches in particular differ (and have to differ) for the boundary contribution proportional to $\delta(1-z)$. (Indeed, the process in ref. [13] generates highly nontrivial twofold integrals $F(m^{2}/\Lambda^{2})$ and $G(m^{2}/\Lambda^{2})$ multiplying $\delta(1-z)$, where $\Lambda$ is the boundary condition of the $X$ plus distributions, with no closed form for $F$ and $G$ away from asymptotic limits for $m$.)
In view of these substantial differences, it is highly nontrivial that the final results in $b_{T}$ space are indeed in full agreement, in particular regarding the latter’s $\delta(1-z)$ boundary condition. In practical terms, we believe that our final momentum-space result in eq. (3.2) compared to the one in eq. (2.30) of ref. [13] nevertheless has significant advantages. A first, practical advantage is that eq. (3.2) is expressed in closed form in terms of elementary functions. A perhaps more important second advantage is that aligning the domains of the plus distributions with the main variables of interest makes it very straightforward to take integral transforms with respect to either $k_{T}$ or $z$, see the discussion above eq. (3.3) and eq. (C.2).
Ref. [13] also considered how the heavy-quark TMD FF factorizes in the bHQET limit $k_{T}\ll m$. At variance with our assumption of $1-z_{H}\sim 1$ in ref. [1] and in eq. (2.17), which amounts to considering wide bins or moments of $z_{H}$, ref. [13] counts $1-z_{H}\sim k_{T}/m\ll 1$ in the bHQET regime, i.e., staying differential in $z_{H}$ but restricting to the peak region, which leads to a TMD bHQET shape function that depends on $z_{H}$, cf. the comment below eq. (2.6) in our ref. [1]. Again, before commenting on the differences between our approaches further, we want to stress that for the power counting of $1-z_{H}\sim k_{T}/m$ assumed by the authors, their factorization of the hadron-level TMD FF in terms of $C_{m}$ and their TMD shape function (which may further be matched onto a collinear shape function, depending on the hierarchy between $\Lambda_{\mathrm{QCD}}$ and $k_{T}$) is entirely correct.
One difference between the two power-counting assumptions on $1-z_{H}$ regards the kind of cross check the respective bHQET calculation provides on the large-mass behavior of the TMD FF. The approach of ref. [13] checks the differential $z$ dependence of $d_{1\,Q/Q}^{(1)}$, but in intermediate steps an identical plus distribution expansion with respect to a very similar, complicated denominator $1/X^{\prime}=1/(k_{T}^{2}+\bar{z}^{2}m^{2})$, which is subject to the same subtleties as the full calculation. In this sense, our calculation provides a strong complementary check of the large-mass behavior since it directly probes the total integral, i.e., a defining property of the distributional result. [^5] Ref. [13] also performed a small-mass consistency check, using the one-loop TMD PDF matching coefficients and a reciprocity relation. This provides a sufficient check at one loop, but reciprocity relations between TMD PDF and FF matching coefficients are known to be subtle at higher orders [47, 46]. Here we stress that the correct all-order relation for the small-mass limit is the one we originally gave in ref. [1], i.e., it has to involve the genuine TMD FF matching coefficients onto twist-2 collinear FFs because the UV matching coefficients are independent of the details of the IR, that is, whether one matches onto a theory at $\mu\sim\Lambda_{\mathrm{QCD}}$ or $\mu\sim m$.
A second and final difference between the power counting assumptions is phenomenological and regards the kinds of differential distributions one may describe in either approach. Our stance here is that while – again – the assumption of ref. [13] is sufficient and entirely correct to describe specific regions of the multi-differential $(z_{H},k_{T})$ distribution, it is not the unique power-counting assumption on $1-z_{H}$, $k_{T}$, $m$ that would yield a leading-power factorization theorem involving bHQET modes, in particular not when different possible scalings of $\Lambda_{\mathrm{QCD}}$ are also considered. As one example, we point to our analysis in section 6, where we continued to count $m\sim k_{T}$, but still took $1-z_{H}$ small, and likewise obtained a leading-power factorization theorem, in that case also involving a purely collinear bHQET shape function, but nontrivial $k_{T}$-dependent correlators of lightlike Wilson lines at the scale $\mu\sim k_{T}$. In light of this, we believe that capturing the full $(z_{H},k_{T})$ dependence will require dedicated future studies combining all of these results, while for the time being and within the scope of exploratory theory studies (or, necessarily less differential, initial experimental measurements) the restriction to wide bins or total integrals over $z_{H}$ is reasonable.
9 Conclusions
In this paper, we calculated all heavy-quark transverse-momentum dependent fragmentation functions (TMD FFs) to next-to-leading order (NLO) in the strong coupling. We provided explicit results in transverse momentum ($k_{T}$) space – both for the differential and the cumulative distribution – and in the conjugate position ($b_{T}$) space, retaining the exact dependence on $k_{T}/m$ and $b_{T}m$, respectively, with $m$ the mass of the heavy quark. Our calculations provide the last missing, key ingredient for a complete next-to-next-to-leading logarithmic (NNLL) description of the transverse momentum distributions of observed heavy hadrons, capturing all quark mass effects. Our results also enable the extension of fixed-order subtraction methods to quasi-collinear limits involving final-state heavy quarks.
We further studied the corresponding TMD fragmentation factors and the TMD light-spin density matrix that arise in boosted Heavy-Quark Effective Theory (bHQET) when the transverse momentum is small compared to the mass. We demonstrated explicitly that they satisfy a two-dimensional set of renormalization group equations (RGEs) with respect to the standard virtuality scale $\mu$ and a dimensionless analog $\rho=\sqrt{\zeta}/m$ of the usual Collins-Soper scale $\zeta$. Here $\rho$ has an immediate physical interpretation as the boost of the heavy hadron in the lab frame. We then proceeded to also evaluate the unpolarized TMD fragmentation factor at NLO, allowing us to cross check our results for the heavy-quark TMD FF in the large-mass and – using known results from the literature – the small-mass limit, finding full agreement. We further showed how to extend the theory of joint threshold and $k_{T}$ resummation to fully account for quark mass effects of $\mathcal{O}(m^{n}b_{T}^{n})$, and used these novel factorization theorems for the heavy-quark TMD FF to obtain a third independent cross check on our NLO result in an asymptotic limit. We also compared our final results for the exact heavy-quark TMD FF in $b_{T}$ space to a recent independent calculation by another group. We found full agreement despite the highly orthogonal ways in which singularities were isolated in the two intermediate momentum-space calculations, which provides a strong independent cross check of this key universal ingredient for heavy-quark TMD predictions. Finally, and as a first application, we leveraged our results to obtain the exact quark-mass dependence of the energy-energy correlator (EEC) in $e^{+}e^{-}$ collisions in the back-to-back limit, where the associated jet functions are given by Mellin moments of our results. As a side note, we pointed out that if the EEC is restricted to heavy hadrons ($\mathrm{EEC}_{H}$), substantial simplifications occur in the back-to-back limit such that the $\mathrm{EEC}_{H}$ at values below the quark mass is given by a simple product of Wilson-line vacuum expectation values, even extending down to nonperturbative values of the observable.
Our results strengthen the theoretical underpinnings of heavy-quark TMD FFs and pave the way towards precision studies of their nonperturbative dynamics, as well as for improving the description of heavy flavor differential distributions in general-purpose Monte-Carlo programs. We look forward to future applications of our results in precision predictions for transverse momentum distributions of heavy quarks and hadrons, and to comparing them to future experimental measurements to shed light on the three-dimensional dynamics of heavy-quark hadronization.
Acknowledgments
We thank Iain Stewart and Kyle Lee for helpful discussions. We thank the Erwin Schrödinger Institute for hospitality while parts of this work were performed. This work was supported in part by the Office of Nuclear Physics of the U.S. Department of Energy under Contract No. DE-SC0011090. Z.S. was also supported by a fellowship from the MIT Department of Physics. R.v.K. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002090 COLORFREE). J.M. was supported by the D-ITP consortium, a program of NWO that is funded by the Dutch Ministry of Education, Culture and Science (OCW).
Appendix A Notation and conventions
A.1 Lightcone coordinates
We decompose four-vectors $p^{\mu}$ in terms of lightlike vectors $n^{\mu}$, ${\bar{n}}^{\mu}$ with $n^{2}=\bar{n}^{2}=0$ and $n\cdot{\bar{n}}=2$,
$p^{\mu}=\bar{n}\cdot p\,\frac{n^{\mu}}{2}+n\cdot p\,\frac{\bar{n}^{\mu}}{2}+p_{\perp}^{\mu}\equiv p^{-}\frac{n^{\mu}}{2}+p^{+}\frac{\bar{n}^{\mu}}{2}+p_{\perp}^{\mu}\equiv(p^{-},p^{+},p_{\perp})\,,$ (A.1)
such that e.g. $p^{2}=p^{-}p^{+}+p_{\perp}^{2}$. We take collinear momenta to have large components $p^{-}\gg p_{\perp}\gg p^{+}$. We always take transverse vectors with subscript $\perp$ to be Minkowskian, $p_{\perp}^{2}\equiv p_{\perp}\cdot p_{\perp}\leq 0$, and denote their magnitude by $p_{T}=\sqrt{-p_{\perp}^{2}}$ such that $p_{T}^{2}\geq 0$ is the square of a real number.
A.2 Plus distributions
One-dimensional plus distributions are defined through their action on test functions $g(x)$,
$\int\!\mathrm{d}x\bigl{[}f(x)\bigr{]}_{+}\,g(x)\equiv\int\!\mathrm{d}x\,f(x)\bigl{[}g(x)-g(0)\,\Theta(1-x)\bigr{]}\,,$ (A.2)
where $f(x)$ diverges at most as $x^{-1-\alpha}$ with $\alpha<1$ for $x\to 1$ and $g(x)$ is assumed to be differentiable as $x\to 0$. Integrals of $[f(x)]_{+}$ against the constant test function over the domain $[0,1]$ vanish by construction. The generalization of eq. (A.2) to two dimensions is given in eq. (3.1) in the main text. For logarithmic plus distributions with homogeneous power counting $\sim 1/x$ and their analogs for dimensionful variables, we further define the shorthands
$\mathcal{L}_{n}(x)\equiv\Bigl{[}\frac{\Theta(x)\,\ln^{n}x}{x}\Bigr{]}_{+}\,,\quad\mathcal{L}_{n}(k,\mu)\equiv\frac{1}{\mu}\mathcal{L}_{n}\Bigl{(}\frac{k}{\mu}\Bigr{)}\,,\quad\mathcal{L}_{n}(k_{T}^{2},\mu^{2})\equiv\frac{1}{\mu^{2}}\mathcal{L}_{n}\Bigl{(}\frac{k_{T}^{2}}{\mu^{2}}\Bigr{)}\,.$ (A.3)
The definition in eq. (A.2) enables expansions in a regulator $\alpha\sim\epsilon,\eta$ as follows,
$\frac{\Theta(x)}{x^{1+\alpha}}$ $=-\frac{\delta(x)}{\alpha}+\Bigl{[}\frac{\Theta(x)}{x^{1+\alpha}}\Bigr{]}_{+}$ $=-\frac{\delta(x)}{\alpha}+\mathcal{L}_{0}(x)-\alpha\mathcal{L}_{1}(x)+\mathcal{O}(\alpha^{2})\,,$ (A.4)
$\frac{\Theta(x)\,\ln x}{x^{1+\alpha}}$ $=-\frac{\delta(x)}{\alpha^{2}}+\Bigl{[}\frac{\Theta(x)\,\ln x}{x^{1+\alpha}}\Bigr{]}_{+}$ $=-\frac{\delta(x)}{\alpha^{2}}+\mathcal{L}_{1}(x)+\mathcal{O}(\alpha)\,.$ (A.5)
To relate plus distributions $\mathcal{L}_{n}(k,\mu)$ with different boundary conditions $\mu_{1,2}$, it is useful to shift one of the boundary conditions by making use of identities like
$\mathcal{L}_{0}(k,\mu_{2})$ $=\mathcal{L}_{0}(k,\mu_{1})+\delta(k)\,\ln\frac{\mu_{1}}{\mu_{2}}\,,$
$\mathcal{L}_{1}(k,\mu_{2})$ $=\mathcal{L}_{1}(k,\mu_{1})+\mathcal{L}_{0}(k,\mu_{1})\,\ln\frac{\mu_{1}}{\mu_{2}}+\delta(k)\,\frac{1}{2}\ln^{2}\frac{\mu_{1}}{\mu_{2}}\,.$ (A.6)
Appendix B Perturbative ingredients
In this appendix we collect various perturbative ingredients from elsewhere in the literature in our notation, as used for cross-checks of our results in the main text. The following expressions are all accurate to one-loop order.
Hard function.
The hard function describing the pair production of quarks through a virtual photon or $Z$ boson in $e^{+}e^{-}$ collisions is given by
$\mathcal{H}_{ee\to ij}(Q^{2},\mu)$ $=\delta_{i\bar{j}}\,\Bigl{\{}e_{i}^{2}-2v_{e}v_{i}e_{i}\operatorname{Re}\bigl{[}P_{Z}(Q^{2})\bigr{]}+(v_{e}^{2}+a_{e}^{2})(v_{i}^{2}+a_{i}^{2})\bigl{\lvert}P_{Z}(Q^{2})\bigr{\rvert}^{2}\Bigr{\}}\bigl{\lvert}C_{q}\bigr{\rvert}^{2}\,,$
$C_{q}(Q^{2},\mu)$ $=1+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-\ln^{2}\frac{-Q^{2}-\mathrm{i}0}{\mu^{2}}+3\ln\frac{-Q^{2}-\mathrm{i}0}{\mu^{2}}-8+\frac{\pi^{2}}{6}\Bigr{]}\,.$ (B.1)
Here we have kept the contribution from $Z$ boson exchange and $Z$-photon interference, as relevant for measurements on the $Z$ pole, where $P_{Z}(Q^{2})=Q^{2}/(Q^{2}-m_{Z}^{2}+\mathrm{i}\Gamma_{Z}m_{Z})$ is the reduced $Z$ propagator and $e_{f}$ ($v_{f}$, $a_{f}$) are the electromagnetic charge (vector, axial coupling to the $Z$) of a fermion $f$.
Large-mass and joint-resummation limits.
The combined collinear and soft mass matching coefficient $C_{m}$, see ref. [1] for our notation, is given by [43]
$C_{m}(m,\mu,\zeta)=1+\frac{\alpha_{s}C_{F}}{4\pi}\biggl{(}4\ln^{2}\frac{\mu}{m}+2\ln\frac{\mu}{m}+4+\frac{\pi^{2}}{6}\biggr{)}+\mathcal{O}\bigl{(}\alpha_{s}^{2}\bigr{)}\,.$ (B.2)
The $\zeta$ dependence arises from the rapidity renormalization of secondary soft heavy-quark contributions starting at $\mathcal{O}(\alpha_{s}^{2})$. The perturbative HQET shape function is given by [5, 6]
$\hat{F}\Bigl{(}\frac{k^{-}}{v^{-}},\mu\Bigr{)}=\delta\Bigl{(}\frac{k^{-}}{v^{-}}\Bigr{)}+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-8\mathcal{L}_{1}\Bigl{(}\frac{k^{-}}{v^{-}},\mu\Bigr{)}-4\mathcal{L}_{0}\Bigl{(}\frac{k^{-}}{v^{-}},\mu\Bigr{)}-\frac{\pi^{2}}{6}\delta\Bigl{(}\frac{k^{-}}{v^{-}}\Bigr{)}\Bigr{]}$ (B.3)
The renormalized TMD soft function for both the $\eta$ regulator of ref. [39, 40] and the exponential regulator of ref. [67] reads
$\sqrt{S}(b_{T},\mu,\nu)=1+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-L_{b}^{2}+4L_{b}\ln\frac{\mu}{\nu}-\frac{\pi^{2}}{6}\Bigr{]}\,.$ (B.4)
The joint-resummation collinear-soft function is given by [68]
$\mathcal{S}(\ell^{-},b_{T},\mu,\nu)=\delta(\ell^{-})+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-8L_{b}\,\mathcal{L}_{0}(\ell^{-},\nu)\Bigr{]}\,.$ (B.5)
Small-mass limit.
The massless TMD FF matching coefficients are given by [73, 29, 47]
$\mathcal{J}_{q/q}(z,b_{T},\mu,\zeta)$ $=\delta(1-z)+\frac{\alpha_{s}C_{F}}{4\pi}\biggl{[}\frac{4(1+z^{2})\ln(z)}{1-z}-{4L_{b}\mathcal{L}_{0}(1-z)}$
$\quad+\delta(1-z)\Bigl{(}-{L_{b}^{2}}-2L_{b}\ln\frac{\zeta}{\mu^{2}}-\frac{\pi^{2}}{6}\Bigr{)}+{2L_{b}(z+1)}+2(1-z)\biggr{]}\,,$
$\mathcal{J}_{g/q}(z,b_{T},\mu,\zeta)$ $=\frac{\alpha_{s}C_{F}}{4\pi}\biggl{[}\frac{4(z^{2}-2z+2)\ln z}{z}-\frac{2L_{b}(z^{2}-2z+2)}{z}+{2z}\biggr{]}\,,$
$\mathcal{J}_{g/g}(z,b_{T},\mu,\zeta)$ $=\delta(1-z)+\frac{\alpha_{s}C_{A}}{4\pi}\biggl{[}\delta(1-z)\Bigl{(}-{L_{b}^{2}}-2L_{b}\ln\frac{\zeta}{\mu^{2}}-\frac{\pi^{2}}{6}\Bigr{)}-{4L_{b}\mathcal{L}_{0}(1-z)}$
$\quad+\frac{8(z^{2}-z+1)^{2}\ln z}{(1-z)z}+\frac{4L_{b}(z^{3}-z^{2}+2z-1)}{z}\biggr{]}$
$\mathcal{J}_{q/g}(z,b_{T},\mu,\zeta)$ $=\frac{\alpha_{s}T_{F}}{4\pi}\bigl{[}2(4z^{2}-4z+2)\ln z-2L_{b}(2z^{2}-2z+1)+4(1-z)z\bigr{]}\,.$ (B.6)
The collinear heavy-quark FFs $d_{Q/Q}(z,\mu)$ and $d_{Q/g}(z,\mu)$ are given by [48]
$d_{Q/Q}(z,\mu)$ $=\delta(1-z)+\frac{\alpha_{s}}{4\pi}C_{F}\Bigl{\{}\ln\frac{\mu^{2}}{m^{2}}\bigl{[}4\mathcal{L}_{0}(1-z)+3\delta(1-z)-2(z+1)\bigr{]}$
$\quad-4\mathcal{L}_{0}(1-z)-8\mathcal{L}_{1}(1-z)+4\delta(1-z)+2(z+1)\bigl{[}2\ln(1-z)+1\bigr{]}\Bigr{\}}\,,$
$d_{Q/g}(z,\mu)=$ $\frac{\alpha_{s}}{4\pi}T_{F}\bigl{[}2z^{2}+2(1-z)^{2}\bigr{]}\ln\frac{\mu^{2}}{m^{2}}\,,\,.$ (B.7)
The decoupling kernels relating collinear FFs in theories with $n_{f}=n_{\ell}+1$ and $n_{\ell}$ active flavors are given by [49]
$\mathcal{M}_{g/g}(z,\mu)$ $=\delta(1-z)+\frac{\alpha_{s}T_{F}}{4\pi}\,\delta(1-z)\Bigl{(}-\frac{4}{3}\Bigr{)}\ln\frac{\mu^{2}}{m^{2}}$
$\mathcal{M}_{g/Q}(z,\mu)$ $=\frac{\alpha_{s}}{4\pi}C_{F}\frac{1+(1-z)^{2}}{z}\Bigl{(}2\ln\frac{\mu^{2}}{m^{2}}-2-4\ln z\Bigr{)}\,.$ (B.8)
In both eqs. (B) and (B) we have rewritten the original results in terms of the minimal distributional basis of $\delta(1-z)$ and $\mathcal{L}_{n}(1-z)$ using eq. (3.37).
Appendix C Results for Mellin moments
We first consider the Mellin moments of the momentum-space TMD FFs and matching coefficients, which are defined in analogy to eq. (7.1) as
$d_{1\,Q/i}^{[N]}(k_{T},\mu,\zeta)$ $\equiv\int\!\mathrm{d}z\,z^{N+2}d_{1\,Q/i}(z,k_{T},\mu,\zeta)\,,$
$\mathcal{J}_{k/i}^{[N]}(k_{T},m,\mu,\zeta)$ $\equiv\int\!\mathrm{d}z\,z^{N}\mathcal{J}_{k/i}(z,k_{T},m,\mu,\zeta)\,.$ (C.1)
For $d_{1\,Q/Q}$, the relevant two-dimensional plus distribution simply turns into
$\int_{0}^{1}\!\mathrm{d}z\,z^{N}\,\bigl{[}f(x,z)\bigr{]}_{+,+}=\Bigl{[}\int_{0}^{1}\!\mathrm{d}z\,\bigl{(}z^{N}-1\bigr{)}f(x,z)\Bigr{]}\,,$ (C.2)
which like the moments of the one-dimensional plus distribution in $z$ can readily be evaluated for general $N$. The remaining one-dimensional plus brackets in eq. (C.2) act on the $x=k_{T}^{2}/m^{2}$ dependence and can further be decomposed in terms of singular $1/x$ terms and regular terms of $\mathcal{O}(x^{0})$ in analogy to eq. (3.37), which is compensated by an additional $\delta(x)$ boundary condition. As an explicit result, we find for $N=1$ at one loop as relevant for section 7:
$d_{1\,Q/Q}^{[1]\,(1)}(k_{T},\mu,\zeta)$ $=\frac{\alpha_{s}C_{F}}{4\pi}\frac{1}{\pi}\biggl{\{}2\ln\frac{\zeta}{\mu^{2}}\mathcal{L}_{0}(k_{T}^{2},\mu^{2})+2\ln\frac{\mu^{2}}{m^{2}}\mathcal{L}_{0}(k_{T}^{2},m^{2})-\ln^{2}\frac{\mu^{2}}{m^{2}}\delta(k_{T}^{2})+3\ln\frac{\mu^{2}}{m^{2}}\delta(k_{T}^{2})$
$\qquad+\frac{1}{m^{2}}\frac{3\pi(-3-18x+x^{2})-\sqrt{x}(-25+9x+45x^{2}+11x^{3})}{\sqrt{x}(1+x)^{4}}$
$\qquad+\frac{1}{m^{2}}\frac{-3\sqrt{x}(1+24x+9x^{2}+2x^{3})\ln x}{\sqrt{x}(1+x)^{4}}-2\mathcal{L}_{0}(k_{T}^{2},m^{2})+\delta(k_{T}^{2})\Bigl{(}4+\frac{\pi^{2}}{6}\Bigr{)}\biggr{\}}\,,$
$d_{1\,Q/g}^{[1]\,(1)}(k_{T},\mu,\zeta)$ $=\frac{\alpha_{s}T_{F}}{4\pi}\frac{1}{\pi m^{2}}\frac{2x(6+x)-30\sqrt{x}\arctan\sqrt{x}-3(-6+x)\ln(1+x)}{3x^{3}}\,,$
$\mathcal{J}_{1\,g/Q}^{[1]\,(1)}(k_{T},m,\mu,\zeta)$ $=\frac{\alpha_{s}C_{F}}{4\pi}\frac{1}{\pi m^{2}}\frac{2(1+4x)}{3(1+x)^{2}}\,.$ (C.3)
Refer to captionRefer to caption
Figure -1074: Left: Numerical Mellin moments $\mathcal{R}^{[N]}(y)$ of the finite remainder term in $d_{1\,Q/Q}$. Right: Numerical Mellin moments $d_{1\,Q/g}^{[N]\,(1)}$ of the gluon-to-massive-quark TMD FF.
Turning to the Mellin moments of the $b_{T}$-space TMD FFs in eq. (7.1), we may use standard Mellin transforms of plus distributions to transform the leading $z\to 1$ terms in $d_{1\,Q/Q}$ in eq. (3.4),
$d_{1\,Q/Q}^{[N]\,(1)}(b_{T},\mu,\zeta)$
$=\frac{\alpha_{s}C_{F}}{4\pi}\bigg{\{}-2L_{b}\ln\frac{\zeta}{m^{2}}+4\ln^{2}\frac{\mu}{m}+6\ln\frac{\mu}{m}-L_{y}^{2}+4-\frac{\pi^{2}}{6}+4(1+L_{y})\bigl{[}\psi(1+N)+\gamma_{E}\bigr{]}$
$\qquad\qquad\quad-8\Bigl{[}-\tfrac{1}{2}\psi^{(1)}(1+N)+\tfrac{1}{2}\bigl{[}\psi(1+N)+\gamma_{E}\bigr{]}^{2}+\frac{\pi^{2}}{12}\Bigr{]}+\tilde{\mathcal{R}}^{[N]}(b_{T}m)\bigg{\}}\,,$ (C.4)
where $\psi(x)$ ($\psi^{(1)}(x)$) is (the first derivative of) the Digamma function. We were not able to find a closed-form expression for the Mellin transform ~$\tilde{\mathcal{R}}^{[N]}(y)\equiv\int_{0}^{1}\!\mathrm{d}z\,z^{N}\,\tilde{\mathcal{R}}(z,y)$ of the finite remainder term, but implementing the $z$ integral of eq. (3.4) directly leads to stable numerical results (shown in the left panel of figure -1074 for reference) that could readily be interpolated. For $\mathcal{J}_{g/Q}$, the Mellin transform for general $N$ is trivial since $z$ does not appear in the arguments of the Bessel $K$ functions,
[…]
For $d_{1\,Q/g}^{[N]\,(1)}$, a closed-form expression in terms of ${}_{1}F_{2}$ hypergeometric and Meijer $G$ functions may be obtained for general $N$ from its definition, but is no more instructive than numerical $z$ moments of eq. (3.5). We show the latter in the right panel of figure -1074 for reference.

### v3

@@
 [16] for recent dedicated projections for polarized collinear parton distribution functions (PDFs) and FFs involving heavy quarks at the EIC.
-However, the applicability of heavy-quark TMD FFs is much wider: Applications of immediate phenomenological interest range from $Z$$+$hadron [17] and dihadron [18] azimuthal decorrelations in $pp$ and $pA$ collisions to hadron-in-jet transverse-momentum distributions [19, 20] and energy-energy correlators in the back-to-back limit in $e^{+}e^{-}$ [21], $ep$ and $eA$ [22, 23], and $pp$ collisions [24, 25].
+However, the applicability of heavy-quark TMD FFs is much wider: Applications of immediate phenomenological interest range from $Z$$+$hadron [17] and dihadron [18] azimuthal decorrelations in $pp$ and $pA$ collisions to hadron-in-jet transverse-momentum distributions [19, 20], transverse momentum-like event shapes (extending the calculation of mass effects in thrust-like event shapes [21, 22, 23, 24]), and energy-energy correlators in the back-to-back limit in $e^{+}e^{-}$ [25], $ep$ and $eA$ [26, 27], and $pp$ collisions [28, 29].
 In all of these cases, factorization formulas involving TMD FFs have been derived, and the NLO results we present in this paper enable one to fully account for the effect of quark masses and accurately capture their highly nontrivial interplay with transverse momenta in all of these processes.
-Figure -1079: The probability density $\mathrm{d}\mathcal{P}_{Q\bar{Q}\to(Qg)\bar{Q}}$ for a gluon to split off a heavy quark in the quasi-collinear limit, which coincides with the heavy-quark TMD FF $d_{1\,Q/Q}$ for finite collinear gluon momentum fraction $\bar{z}>0$ and transverse momentum $k_{T}>0$.
+Figure -1199: The probability density $\mathrm{d}\mathcal{P}_{Q\bar{Q}\to(Qg)\bar{Q}}$ for a gluon to split off a heavy quark in the quasi-collinear limit, which coincides with the heavy-quark TMD FF $d_{1\,Q/Q}$ up to a factor of $z^{2}$ for finite collinear gluon momentum fraction $\bar{z}>0$ and transverse momentum $k_{T}>0$.
 On the rear walls of the plot we show the integrals of $d_{1\,Q/Q}$ over one of $k_{T}$ or $z$ as predicted by the NLO calculation we perform in this paper, showcasing that the TMD regularization and renormalization procedure renders them finite despite the intricate singular behavior of the integrand.
-The same is true for the total integral over both $z$ and $k_{T}$, which is not shown in the plot, but at the values of the scales chosen here is equal to $\frac{\alpha_{s}C_{F}}{4\pi}\bigl{(}5-\frac{\pi}{2}-\frac{2\pi^{2}}{3}\bigr{)}$.
+The same is true for the total integral over both $z$ and $k_{T}<m$, which is not shown in the plot, but at the values of the scales chosen here is equal to $\frac{\alpha_{s}C_{F}}{4\pi}\bigl{(}5-\frac{\pi}{2}-\frac{2\pi^{2}}{3}\bigr{)}\approx-0.067$.
 For numerical illustration, we take $\alpha_{s}(\mu=m)=0.2$.
-The heavy-quark TMD FF is intimately connected with the universal gluon splitting probability off a heavy-quark in the quasi-collinear limit [26, 27], e.g.
+The heavy-quark TMD FF is intimately connected with the universal gluon splitting probability off a heavy quark in the quasi-collinear limit [30, 31], e.g.
 after heavy-quark pair production has taken place,
@@
 where $m$ is the heavy-quark mass, $k_{T}$ is the transverse momentum of the gluon with respect to the heavy quark, and $z$ is the collinear momentum fraction retained by the heavy quark.
-The quasi-collinear splitting amplitude is shown in a three-dimensional plot as a function of $k_{T}$ and the momentum fraction $\bar{z}\equiv 1-z$ carried by the gluon in figure -1079.
+The quasi-collinear splitting amplitude is shown in a three-dimensional plot as a function of $k_{T}$ and the momentum fraction $\bar{z}\equiv 1-z$ carried by the gluon in figure -1199.
 Famously, and in contrast to the behavior of light partons, the splitting amplitude is suppressed for energetic gluons with large momentum fractions $\bar{z}\gtrsim 0.2$, an effect known as the dead cone [32].
@@
 (1.1) to $\bar{z}\to 0$ and/or $k_{T}\to 0$ in a way that is manifestly finite in four spacetime dimensions.
-As a demonstration, we include on the rear walls of the three-dimensional plot in figure -1079 the integral of the heavy-quark TMD FF over one of $z$ or $k_{T}\leq m$, respectively, as predicted by our calculations in this paper, showcasing that they are indeed individually finite.
+As a demonstration, we include on the rear walls of the three-dimensional plot in figure -1199 the integral of the heavy-quark TMD FF over one of $z$ or $k_{T}\leq m$, respectively, as predicted by our calculations in this paper, showcasing that they are indeed individually finite.
 Similarly, the total integral over both variables $0\leq z\leq 1$ and $0\leq k_{T}\leq m$ (and thus any integral over a subset of the plane) is also finite in the usual distributional sense, and for the indicated values of the scales simply evaluates to $\frac{\alpha_{s}C_{F}}{4\pi}\bigl{(}5-\frac{\pi}{2}-\frac{2\pi^{2}}{3}\bigr{)}$.
@@
 In ref.
-[1] we showed that up to a factor of the total probability for $Q$ to fragment into $H$, the distribution differential in $k_{T}=P_{X,T}\sim M_{H}$ and the lightcone momentum fraction $z_{H}$ of the original quark retained by $H$ is governed by a new perturbative matching coefficient, the partonic heavy-quark TMD FF $d_{1\,Q/Q}(z,k_{T},\mu,\zeta)$,
+[1] we showed that up to a factor of the total probability $\chi_{H}$ for $Q$ to fragment into $H$, the distribution differential in $k_{T}=P_{X,T}\sim M_{H}$ and the lightcone momentum fraction $z_{H}$ of the original quark retained by $H$ is governed by a new perturbative matching coefficient, the partonic heavy-quark TMD FF $d_{1\,Q/Q}(z,k_{T},\mu,\zeta)$,
 $D_{1\,H/Q}(z_{H},b_{T},\mu,\zeta)=d_{1\,Q/Q}(z_{H},b_{T},\mu,\zeta)\,\chi_{H}+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{m}\Bigr{)}+\mathcal{O}(\Lambda_{\mathrm{QCD}}b_{T})\,.$ (2.1)
@@
 $d_{1\,Q/Q}(z,b_{T},\mu,\zeta)$ $=\lim_{\epsilon\to 0}Z_{\mathrm{UV}}(\mu,\zeta,\epsilon)\lim_{\eta\to 0}\Bigl{[}d_{1\,Q/Q}^{\mathrm{bare}}(z,b_{T},\epsilon,\eta,\zeta/\nu^{2})\,\sqrt{S}(b_{T},m,\epsilon,\eta,\nu)\Bigr{]}\,,$ (2.14)
-where $S$ is the universal bare TMD soft function for the $\eta$ regulator, which cancels all poles of $\eta$ and the associated $\nu$ dependence within the second line and is independent of the heavy-quark mass up to secondary quark mass effects starting at two loops.
+where $S$ is the universal bare TMD soft function for the $\eta$ regulator, which cancels all poles of $\eta$ and the associated $\nu$ dependence, and is independent of the heavy-quark mass up to secondary quark mass effects starting at two loops.
 By contrast, the $\overline{\text{MS}}$ renormalization factor $Z_{\mathrm{UV}}$ for quark TMDs is independent of the mass to all orders by RG consistency with the hard matching coefficient at the hard scattering energy $Q\sim p^{-}/z\gg m\sim k_{T}\sim 1/b_{T}$.
@@
 where $C_{m}(m,\mu,\zeta)$ is generated by separately matching collinear and soft modes at the scale $\mu\sim m$ onto HQET and QCD with $n_{\ell}$ flavors, respectively [47]; explicit expressions for $C_{m}$ and other perturbative ingredients to the order required for our perturbative checks are given in appendix B.
+We note that eq.
+(2.17) may appear counterintuitive, given that the prediction of HQET in eq.
+(2.18) – after assuming this counting – comes out proportional to $\delta(1-z_{H})$.
+The proper way to understand eq.
+(2.17) in this context is that it may be satisfied not just point by point in $z_{H}$ far from the endpoint (where HQET predicts that the heavy-quark TMD FF is suppressed), but in particular is also satisfied when integrating over a wide bin in $z_{H}$ from $1$ down to some bin boundary $z_{H}^{\mathrm{cut}}$ with $1-z_{H}^{\mathrm{cut}}\sim 1$, or when taking Mellin moments $z_{H}^{N}$ with $N\ll m/\Lambda_{\mathrm{QCD}}$ not large.
+In these latter two cases, the prediction from HQET proportional to $\delta(1-z_{H})$ can be easily understood as all bin integrals and Mellin moments becoming equal and independent of $z_{H}^{\mathrm{cut}}$ and $N$ up to corrections of $\mathcal{O}(1/m)$ [1].
+We point out that the assumption in eq.
+(2.17) in fact also underlies the “full theory” factorization for $k_{T}\sim m$ in eq.
+(2.1).
+Lowering $1-z_{H}\ll 1$ at fixed $k_{T}\sim m$ in that case requires further resummation and – in the extreme limit – a different factorization of nonperturbative physics, which we present in Section 6.
 In eq.
@@
 where $C_{1}$ will be the target of our calculation in section 4.3.
-Here we have also introduced the shorthand $\rho\equiv\bar{n}\cdot v=\sqrt{\zeta}/m$ for the boost of the hadron; in ref.
+Here we have also introduced the shorthand $\rho\equiv v^{-}=\sqrt{\zeta}/m$ for the boost of the hadron; in ref.
 [1] we argued from consistency with TMD factorization that $\chi_{1,H}$ develops an anomalous dependence on this variable governed by the Collins-Soper kernel, and we will make this expectation more explicit in section 4.1.
@@
 $D_{H/i}^{(n_{\ell}+1)}(z_{H},\mu)$ $=d_{Q/i}(z_{H},\mu)\,\chi_{H}+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{m}\Bigr{)}\,,$
-$D_{h/i}^{(n_{\ell}+1)}(z_{h},\mu)$ $=\sum_{j}\int\!\frac{\mathrm{d}z}{z}\,\mathcal{M}_{j/i,T}(z,m,\mu)\,D_{h/j}^{(n)}\Bigl{(}\frac{z_{h}}{z},\mu\Bigr{)}\,.$ (2.23)
+$D_{h/i}^{(n_{\ell}+1)}(z_{h},\mu)$ $=\sum_{j}\int\!\frac{\mathrm{d}z}{z}\,\mathcal{M}_{j/i,T}(z,m,\mu)\,D_{h/j}^{(n_{\ell})}\Bigl{(}\frac{z_{h}}{z},\mu\Bigr{)}\,.$ (2.23)
 Here $d_{Q/i}(z_{H},\mu)$ is the partonic collinear heavy-quark FF and $\mathcal{M}_{j/i,T}$ is the timelike matching function governing the flavor decoupling in light-hadron collinear FFs.
@@
 […]
-The two contributing real-emission diagrams in Feynman gauge are given in figure -1078 (a) and (b), where $X=g$ is a single gluon.
+The two contributing real-emission diagrams in Feynman gauge are given in figure -1198 (a) and (b), where $X=g$ is a single gluon.
 Including the mirror diagram for (a), their contributions to $d_{1\,Q/Q}^{\mathrm{bare}\,(1)}$ evaluate to
@@
 This expansion is valid here because the remaining terms in eq.
-(3.11) are still dimensionally regulated at $\bar{z}=0$ as $k_{T}\to 0$ i.e., the remainder of the expression is a good test function for general $\epsilon$.
+(3.11) are still dimensionally regulated at $\bar{z}=0$ as $k_{T}\to 0$ i.e., the remainder of the expression is a good test function for finite $\epsilon\neq 0$.
 Since the mass enters diagram (a) only through the combination $m\bar{z}$, the action of the $\delta(\bar{z})/\eta$ term on the rest of the diagram is to reduce it to the massless limit.
@@
 where $f(x,\epsilon)\sim x^{-1-\epsilon}+\mathcal{O}(x^{0})$ has a dimensionally regulated singular limit $x\to 0$ on the first left-hand side.
-By contrast, on the right-hand side $f(x,\epsilon)$ may be expanded in $\epsilon$ since each term in the expansion is now dimensionally regulated, whereas the explicit poles are typically isolated in the Laurent series of the total integral $F(\epsilon)$.
+By contrast, on the right-hand side $f(x,\epsilon)$ may be expanded in $\epsilon$ since each term in the expansion is now regulated by the plus prescription, whereas the explicit poles are typically isolated in the Laurent series of the total integral $F(\epsilon)$.
 Of course, this is precisely how relations like eq.
@@
 (3.16), such that $d_{1\,Q/Q}^{\mathrm{bare}\,(a+s)}$ is finite as $\epsilon\to 0$ for $z<1$ for any $x\geq 0$.
-The double pole $\tfrac{1}{\epsilon}\delta(\bar{z})\delta(k_{T}^{2})$ similarly cancels.
-The second diagram in figure -1078 does not require rapidity regularization. In order to expand it in the dimensional regulator, we repeat the above procedure, where we define
+The double pole $\tfrac{1}{\epsilon^{2}}\delta(\bar{z})\delta(k_{T}^{2})$ similarly cancels.
+The second diagram in figure -1198 does not require rapidity regularization. In order to expand it in the dimensional regulator, we repeat the above procedure, where we define
 $d_{1\,Q/Q}^{\mathrm{bare}\,(b)}$ $\equiv\frac{\alpha_{s}C_{F}}{4\pi}\frac{e^{\epsilon\gamma_{E}}}{\Gamma(1-\epsilon)}\frac{1}{\pi z^{2-2\epsilon}}\frac{\mu^{2\epsilon}}{m^{2+2\epsilon}}\,f^{(b)}(k_{T}^{2}/m^{2},z,\epsilon)\,,$
@@
 3.2 Virtual contributions and UV renormalization
-The virtual diagram (c) from figure -1078 is common to any collinear matrix element involving massive fermions and can be evaluated in a straightforward way using integration by contours for the $\ell^{+}$ component of the loop momentum $\ell$, Feynman parametrization for the $\ell_{\perp}$ integral, and an analytic $\ell^{-}$ integral at the end.
+The virtual diagram (c) from figure -1198 is common to any collinear matrix element involving massive fermions and can be evaluated in a straightforward way using integration by contours for the $\ell^{+}$ component of the loop momentum $\ell$, Feynman parametrization for the $\ell_{\perp}$ integral, and an analytic $\ell^{-}$ integral at the end.
 Including the mirror diagram, the result reads
@@
 (3.19), i.e., as a two-dimensional plus distribution with nontrivial functions of $z$ and $k_{T}$ living on the respective boundaries and an overall boundary contribution at $z=1$ and $k_{T}=0$.
-This structure was already illustrated in figure -1079, where the partial integrals over one of the variables can immediately be read off from the respective boundary terms.
+This structure was already illustrated in figure -1199, where the partial integrals over one of the variables can immediately be read off from the respective boundary terms.
 Of course, the total contribution in the bulk at $z<1$ and $k_{T}>0$, which can simply be read off from the content of the $[\dots]_{+,+}$ brackets accounting for the prefactor of $1/z^{2}$, is still equal to the simple result we found in eq.
@@
 3.5 Nonvalence channels at $\mathcal{O}(\alpha_{s})$
-Figure -1077: Diagrams contributing to the gluon-initiated partonic heavy-quark TMD FF and the mass-dependent TMD FF matching coefficients at $\mathcal{O}(\alpha_{s})$.
+Figure -1197: Diagrams contributing to the gluon-initiated partonic heavy-quark TMD FF and the mass-dependent TMD FF matching coefficients at $\mathcal{O}(\alpha_{s})$.
 Heavy quark propagators and field insertions are indicated in red.
@@
 For completeness we also compute the partonic TMD FFs (or TMD FF matching coefficients onto collinear distributions) for all remaining partonic channels involving heavy quarks.
-The relevant diagrams at $\mathcal{O}(\alpha_{s})$ are given in figure -1077, where we labeled the identified final-state “hadron” by the momentum $p$.
-Note that for diagrams (b) and (c) compared to the otherwise identical diagrams in figure -1078, this also changes the definition of $k_{\perp}$, as the coordinate system needs to be adjusted such that the gluon now has vanishing transverse momentum.
+The relevant diagrams at $\mathcal{O}(\alpha_{s})$ are given in figure -1197, where we labeled the identified final-state “hadron” by the momentum $p$.
+Note that for diagrams (b) and (c) compared to the otherwise identical diagrams in figure -1198, this also changes the definition of $k_{\perp}$, as the coordinate system needs to be adjusted such that the gluon now has vanishing transverse momentum.
 Results for antiquarks are identical at this order.
@@
 This is the same soft function as appears in light-quark TMD PDFs or FFs at transverse momenta below the heavy-quark threshold: The fact that it also arises here and cancels the rapidity divergences as $\eta\to 0$ in the bare light-spin density matrix follows from consistency of the TMD factorization for flavor-changing hard processes, e.g.
-charged-current DIS $e^{-}d\to\nu_{e}\,c$111 We assume as in section 2.1 that a rapidity regulator $\eta$ is used for which soft (zero-bin) subtractions in the bare collinear matrix element are scaleless.
+charged-current DIS $e^{-}\bar{s}\to\nu_{e}\,\bar{c}$.111 We assume as in section 2.1 that a rapidity regulator $\eta$ is used for which soft (zero-bin) subtractions in the bare collinear matrix element are scaleless.
 If this is not the case, it is crucial that zero-bin subtractions are defined by modes that are soft in a frame where the heavy quark is highly boosted, i.e., their interactions with the heavy-quark sector are encoded in lightlike Wilson lines and the multiplicative zero-bin subtraction factor is obtained by taking $\bar{n}\cdot v\to\infty$.
@@
 By contrast, in physical cross sections, the boost is fixed to $\rho=Q/m\gg 1$, i.e., the lab frame in the $e^{+}e^{-}$ case, or the Breit frame for SIDIS.
-The $\rho$-evolution between those scales resums the associated large rapidity logarithms.
+The $\rho$ evolution between those scales resums the associated large rapidity logarithms.
 As in the relativistic case, the two-dimensional evolution equations satisfy an integrability condition,
@@
 4.3 One-loop calculation
-Figure -1076: Left: Diagrams contributing to the total bHQET TMD fragmentation factor in the perturbative domain.
+Figure -1196: Left: Diagrams contributing to the total bHQET TMD fragmentation factor in the perturbative domain.
 Wilson lines along the lightlike ${\bar{n}}$ direction are indicated by the slanted double lines.
@@
 Virtual diagrams with $X$ the vacuum state are scaleless.
-The two contributing real emission diagrams at $\mathcal{O}(\alpha_{s})$ in Feynman gauge are shown in figure -1076b.
+The two contributing real emission diagrams at $\mathcal{O}(\alpha_{s})$ in Feynman gauge are shown in figure -1196b.
 The one-gluon Feynman rule for the $Y_{v}$ Wilson line can e.g.
@@
 $\int\!\mathrm{d}y\,\mathcal{A}^{(a)}(y,\eta)=-\frac{2\pi}{\sin(\eta\pi/2)}=-\frac{4}{\eta}+\mathcal{O}(\eta)\,,\qquad\int\!\mathrm{d}y\,\mathcal{A}^{(b)}(y)=-2\,.$ (4.28)
-The origin of the rapidity divergence, like the structure of the integrands as a whole, has a rather intuitive interpretation, which we illustrate in the right panel of figure -1076b: Emissions at large positive gluon rapidity $y$ in the rest frame of the heavy quark are strongly suppressed for either contributing diagram, which is the well-known dead-cone effect [28].
+The origin of the rapidity divergence, like the structure of the integrands as a whole, has a rather intuitive interpretation, which we illustrate in the right panel of figure -1196b: Emissions at large positive gluon rapidity $y$ in the rest frame of the heavy quark are strongly suppressed for either contributing diagram, which is the well-known dead-cone effect [32].
 For diagram (b), where the static heavy quark $Y_{v}$ both emits and absorbs the gluon, the contribution to the emission probability (dot-dashed blue) peaks at $y=0$ and is suppressed also at large negative rapidities because the contribution must be even in the rest frame.
 In contrast to this, the contribution from diagram (a), where the emission couples to the $W_{\eta}$ that arises from integrating out the interactions with the anticollinear and soft sectors and is required for gauge invariance, tends to a large constant as $y\to-\infty$ (dashed red), leading to a rapidity divergence.
-Of course, this precisely corresponds to the central soft momentum region, and thus after introducing the regulator $e^{-\eta y}$ with $\eta<0$ to dampen the behavior as $y\to\infty$ (solid red), the associated pole can cancel against the soft function as usual.
+Of course, this precisely corresponds to the central soft momentum region, and thus after introducing the regulator $e^{-\eta y}$ with $\eta<0$ to dampen the behavior as $y\to\infty$ (solid red), the associated pole cancels against the soft function as usual.
 Adding the one-loop soft function in eq.
@@
 6 Joint threshold and $q_{T}$ resummation for massive quark fragmentation
-Figure -1075: Relevant modes contributing to the heavy-quark TMD FF as the collinear momentum fraction $z_{H}\to 1$ while counting $m\sim k_{T}$. See the text for a detailed discussion.
+Figure -1195: Relevant modes contributing to the heavy-quark TMD FF as the collinear momentum fraction $z_{H}\to 1$ while counting $m\sim k_{T}$. See the text for a detailed discussion.
 The remarkable simplicity of our final $b_{T}$-space result in eq.
@@
 We stress that we continue to count $m\sim k_{T}\sim 1/b_{T}$.
-In figure -1075 we illustrate the contributing modes in the EFT for this scenario.
-Figure -1075 (a) shows the original, unmodified mode setup that led to eq.
+In figure -1195 we illustrate the contributing modes in the EFT for this scenario.
+Figure -1195 (a) shows the original, unmodified mode setup that led to eq.
 (2.1).
-Here the partonic heavy-quark TMD FF arises from matching collinear modes (green), accounting for the usual contribution of the central soft modes (orange), onto the nonperturbative modes (gray) describing the brown muck, which is isotropic in the rest frame of the heavy quark indicated by the dashed black line.
+Here the partonic heavy-quark TMD FF arises from matching collinear modes (green), accounting for the usual contribution of the central soft modes (orange), onto the nonperturbative modes (gray) describing the formation of the brown muck, which is isotropic in the rest frame of the heavy quark indicated by the dashed black line.
 The collinear modes are kinematically unconstrained in the longitudinal direction because $\bar{z}_{H}\sim 1$.
 Once the modified assumption eq.
-(6.1) is made, collinear modes at the scale $\mu\sim k_{T}\sim m$ can no longer be radiated into the final state.
-Their effect becomes purely virtual and is encoded in the matching coefficient $C_{m}$ (green cross) from matching onto bHQET modes at yet lower virtuality.
+(6.1) is made, energetic collinear modes at the scale $\mu\sim k_{T}\sim m$ can no longer be radiated into the final state, see figure -1195 (b) and (c).
+Their effect becomes purely virtual and is encoded in a matching coefficient $H_{m,n}^{Q\to h_{v}}$ (green cross) from matching onto bHQET modes at lower virtuality.
 Real emissions into the $n$-collinear direction at the scale $\mu\sim k_{T}\sim m$ can instead only have collinear-soft (pink) scaling
@@
 which is the characteristic mode of double-differential $z\to 1$ threshold and $q_{T}$ resummation, often simply dubbed joint resummation [64, 65, 66, 67, 68, 69, 70, 71, 72].222 We note that while ref.
-[64] also dealt with final-state heavy quarks, that work considered the joint resummation for the total transverse momentum of a heavy quark pair with respect to the beam direction, which is a problem that is distinct from the one we consider here.
-The physical picture at scales $\mu\ll k_{T}\sim m$ in this scenario, which is governed by one or several bHQET modes at their characteristic lab-frame rapidity $p^{-}/p^{+}\sim Q^{2}/m$, further depends on the relative size of $\Lambda_{\mathrm{QCD}}/m$ and $\bar{z}_{H}$, specifically whether $\Lambda_{\mathrm{QCD}}/m\sim\bar{z}_{H}$ or $\Lambda_{\mathrm{QCD}}/m\ll\bar{z}_{H}$.
-(Since $\Lambda_{\mathrm{QCD}}$ introduces a fourth scale into the problem, i.e., a third independent dimensionless power counting parameter, either case may hold in general and independent of $k_{T}/Q$ and $m/Q$.) The relevant modes for the case $\Lambda_{\mathrm{QCD}}/m\sim\bar{z}_{H}$ are illustrated in figure -1075 (c), and are in fact closely analogous to the well-understood case of single-inclusive heavy-hadron production near the endpoint in $e^{+}e^{-}$ collisions.
+[68] also dealt with final-state heavy quarks, that work considered the joint resummation for the total transverse momentum of a heavy quark pair with respect to the beam direction, which is a distinct problem from the one we consider here.
+The physical picture at scales $\mu\ll k_{T}\sim m$ in this scenario, which is governed by one or several bHQET modes at their characteristic lab-frame rapidity $p^{-}/p^{+}\sim Q^{2}/m^{2}$, further depends on the relative size of $\Lambda_{\mathrm{QCD}}/m$ and $\bar{z}_{H}$, specifically whether $\Lambda_{\mathrm{QCD}}/m\sim\bar{z}_{H}$ or $\Lambda_{\mathrm{QCD}}/m\ll\bar{z}_{H}$.
+(Since $\Lambda_{\mathrm{QCD}}$ introduces a fourth scale into the problem, i.e., a third independent dimensionless power counting parameter, either case may hold in general and independent of $k_{T}/Q$ and $m/Q$.) The relevant modes for the case $\Lambda_{\mathrm{QCD}}/m\sim\bar{z}_{H}$ are illustrated in figure -1195 (c), and are in fact analogous to the well-understood case of single-inclusive heavy-hadron production near the endpoint in $e^{+}e^{-}$ collisions [5, 6].
 Specifically, the nonperturbative fragmentation process at the scale $\mu\sim\Lambda_{\mathrm{QCD}}\sim\bar{z}_{H}m$ provides a leading contribution to the measured longitudinal hadron momentum in this case (but not to its transverse momentum, where the contribution is power suppressed!).
-This effect is encoded in the well-known collinear bHQET fragmentation shape function $F_{H}(k^{-},\mu)$ [5, 6].
+This effect is encoded in the well-known collinear bHQET fragmentation shape function $F_{H}(k^{-}/v^{-},\mu)$ [5, 6].
 [^3] We use the symbol $F_{H}$ here to distinguish the shape function from the TMD soft function; in ref.
@@
 In total, the factorization of the hadron-level heavy-quark TMD FF in this regime reads, with $\zeta=\omega^{2}$ the Collins-Soper scale,
-$D_{1\,H/Q}(z_{H},b_{T},\mu,\omega^{2})$ $=C_{m}(m,\mu,\omega^{2})\,\int\!\mathrm{d}\ell^{-}\,\mathcal{S}\Bigl{(}b_{T},m,\mu,\frac{\ell^{-}}{\nu}\Bigr{)}\,\sqrt{S}(b_{T},m,\mu,\nu)$
-×$\quad\times\int\!\frac{\mathrm{d}k^{-}}{\bar{n}\cdot v}\,F_{H}\Bigl{(}\frac{k^{-}}{\bar{n}\cdot v},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}_{H}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}$
+$D_{1\,H/Q}(z_{H},b_{T},\mu,\omega^{2})$ $=H_{m,n}^{Q\to h_{v}}\Bigl{(}m,\mu,\frac{\nu}{\omega}\Bigr{)}\,\int\!\mathrm{d}\ell^{-}\,\mathcal{S}(\ell^{-},b_{T},m,\mu,\nu)\,\sqrt{S}(b_{T},m,\mu,\nu)$
+×$\quad\times\int\!\frac{\mathrm{d}k^{-}}{v^{-}}\,F_{H}\Bigl{(}\frac{k^{-}}{v^{-}},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}_{H}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}$
 ×$\quad\times\biggl{[}1+\mathcal{O}(\bar{z}_{H})+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{k_{T}}\Bigr{)}+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{m}\Bigr{)}\biggr{]}\,,$ (6.3)
-where $\mathcal{S}$ and $\sqrt{S}$ are the separately renormalized collinear-soft function from joint resummation [68] and the square root of the standard TMD soft function, respectively.
+where $\mathcal{S}$ and $\sqrt{S}$ are the separately rapidity-renormalized collinear-soft function from joint resummation [72] and the square root of the standard TMD soft function, respectively.
+Similarly, $H_{m,n}^{Q\to h_{v}}$ (defined in ref.
+[47] without the superscript, which we here include for clarity) is the rapidity-renormalized matching coefficient encoding the virtual contributions of massive collinear quark modes $Q$ with label momentum $\omega$ matched onto bHQET fields $h_{v}$.
+As in ref.
+[47], the effect of these modes is purely virtual due to the kinematic constraint $z_{H}\to 1$.
+We note that $H_{m,n}^{Q\to h_{v}}$ also includes the effect of secondary heavy quark loops not directly connected to the operator starting at $\mathcal{O}(\alpha_{s}^{2})$, which are responsible for its rapidity renormalization.
 By evolving all terms in eq.
 (6) from their canonical scales (e.g.
-in Mellin space $N\leftrightarrow z_{H}$ where $\nu_{\mathcal{S}}\sim\bar{n}\cdot v\,\mu_{F_{H}}\sim Q/N$ and the evolution is multiplicative) to the common scale $\mu\sim\sqrt{\zeta}\sim Q$ with the hard function, all large threshold and transverse momentum logarithms present in the heavy-hadron TMD cross section in this regime can be resummed simultaneously.
-It is instructive to compare the above to the refactorization of the light-parton TMD FF in the joint resummation regime $k_{T}\ll\bar{z}_{h}Q\ll Q$, which follows from crossing the light-parton TMD PDF refactorization relations [68],
-$D_{1\,h/q}(z_{h},b_{T},\mu,\omega^{2})$ $=\int\!\mathrm{d}\ell^{-}\,\mathcal{S}\Bigl{(}b_{T},m,\mu,\frac{\ell^{-}}{\nu}\Bigr{)}\,\sqrt{S}(b_{T},m,\mu,\nu)$
+in Mellin space $N\leftrightarrow z_{H}$, where $\nu_{\mathcal{S}}\sim\mu_{F_{H}}v^{-}\sim\omega/N$ and the evolution is multiplicative) to the common scale $\mu\sim\nu\sim Q$ with the hard function, all large threshold and transverse momentum logarithms present in the heavy-hadron TMD cross section in this regime can be resummed simultaneously.
+It is instructive to compare the above to the refactorization of the light-parton TMD FF in the joint resummation regime $k_{T}\ll\bar{z}_{h}Q\ll Q$, which follows from crossing the light-parton TMD PDF refactorization relations [72], accounting for the additional effect of secondary heavy quarks starting at two loops,
+$D_{1\,h/q}(z_{h},b_{T},\mu,\omega^{2})$ $=H_{m,n}^{q\to q}\Bigl{(}m,\mu,\frac{\nu}{\omega}\Bigr{)}\int\!\mathrm{d}\ell^{-}\,\mathcal{S}(\ell^{-},b_{T},m,\mu,\nu)\,\sqrt{S}(b_{T},m,\mu,\nu)$
 ×$\quad\times\int\!\mathrm{d}k^{-}\,D^{\mathrm{thr}}_{h/q}\Bigl{(}\frac{k^{-}}{P_{h}^{-}},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}_{h}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}$
 ×$\quad\times\biggl{[}1+\mathcal{O}(\bar{z}_{H})+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{k_{T}}\Bigr{)}+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{m}\Bigr{)}\biggr{]}\,,$ (6.4)
-where $D^{\mathrm{thr}}_{h/q}$ is the collinear (threshold) fragmentation function, whose renormalization is given by the timelike DGLAP kernels expanded to leading power in $1-z$.
-The product $C_{m}\,F_{H}$ in eq.
-(6) is renormalized like a threshold FF [6], thus eq.
-(6) is indeed RGE consistent by comparison with eq.
-(6).
-We note that both eqs.
-(6) and (6) generically feature a dependence on the heavy quark mass in the soft and collinear-soft function since $mb_{T}\sim 1$, to which we return towards the end of this section.
+[…]
 To make contact with the partonic heavy-quark TMD FF, we next take $\Lambda_{\mathrm{QCD}}\to 0$, which in particular means that now $\Lambda_{\mathrm{QCD}}/m\ll\bar{z}_{H}$.
-The mode setup for this regime is illustrated in figure -1075 (b).
-The only change compared to figure -1075 (c) is that the fragmentation shape function now factorizes into the perturbative bHQET shape function $\hat{F}$ and the total fragmentation probability $\chi_{H}$ [6],
-$F_{H}(k^{-},\mu)=\hat{F}(k^{-},\mu)\,\chi_{H}+\mathcal{O}(\Lambda_{\mathrm{QCD}})\,.$ (6.5)
+The mode setup for this regime is illustrated in figure -1195 (b).
+The only change compared to figure -1195 (c) is that the fragmentation shape function now factorizes into the perturbative bHQET shape function $\hat{F}$ and the total fragmentation probability $\chi_{H}$ [6],
+$F_{H}\Bigl{(}\frac{k^{-}}{v^{-}},\mu\Bigr{)}=\hat{F}\Bigl{(}\frac{k^{-}}{v^{-}},\mu\Bigr{)}\,\chi_{H}+\mathcal{O}(\Lambda_{\mathrm{QCD}})\,.$ (6.5)
 This implies that the joint resummation of the partonic heavy-quark TMD FF is governed by the following refactorization relation,
-$d_{1\,Q/Q}(z_{H},b_{T},\mu,\omega^{2})$ $=C_{m}(m,\mu,\omega^{2})\,\int\!\mathrm{d}\ell^{-}\,\mathcal{S}\Bigl{(}b_{T},m,\mu,\frac{\ell^{-}}{\nu}\Bigr{)}\,\sqrt{S}(b_{T},m,\mu,\nu)$
-×$\quad\times\int\!\frac{\mathrm{d}k^{-}}{\bar{n}\cdot v}\,\hat{F}\Bigl{(}\frac{k^{-}}{\bar{n}\cdot v},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}_{H}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}+\mathcal{O}\bigl{[}(1-z_{H})^{0}\bigr{]}\,.$ (6.6)
+$d_{1\,Q/Q}(z,b_{T},\mu,\omega^{2})$ $=H_{m,n}^{Q\to h_{v}}\Bigl{(}m,\mu,\frac{\nu}{\omega}\Bigr{)}\,\int\!\mathrm{d}\ell^{-}\,\mathcal{S}(\ell^{-},b_{T},m,\mu,\nu)\,\sqrt{S}(b_{T},m,\mu,\nu)$
+×$\quad\times\int\!\frac{\mathrm{d}k^{-}}{v^{-}}\,\hat{F}\Bigl{(}\frac{k^{-}}{v^{-}},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}+\mathcal{O}\bigl{[}(1-z)^{0}\bigr{]}\,.$ (6.6)
 This is precisely the relation governing the leading $1/\bar{z}$ terms in the partonic heavy-quark TMD FF that we were seeking out, and it is straightforward to verify eq.
 (6) at one-loop order against eq.
-(3.4) using the fact that ~$\tilde{\mathcal{R}}(z,y)=\mathcal{O}[(1-z)^{0}]$, the fact that $\bar{n}\cdot v=\omega/m$, and the ingredients given in appendix B.
-In particular, since the soft and soft-collinear function are defined as Wilson line correlators, they can only acquire mass dependence through secondary mass effects starting at $\mathcal{O}(\alpha_{s}^{2}C_{F}T_{F})$.
+(3.4) using the fact that ~$\tilde{\mathcal{R}}(z,y)=\mathcal{O}[(1-z)^{0}]$, the fact that $v^{-}=\omega/m$, and the ingredients given in appendix B.
+In particular, since the soft and collinear-soft function are defined as Wilson line correlators, they can only acquire mass dependence through secondary mass effects starting at $\mathcal{O}(\alpha_{s}^{2}C_{F}T_{F})$.
 In turn, this implies that the mass dependence at $\mathcal{O}(\alpha_{s})$ arises only from that of $C_{m}$, which is logarithmic to all orders, explaining why eq.
 (3.4) has to have the simple logarithmic mass dependence that we found for $\bar{z}\to 1$.
-We stress that starting at $\mathcal{O}(\alpha_{s}^{2})$, the soft and soft-collinear functions do develop a highly nontrivial dependence on $mb_{T}$, which is e.g.
-inherited by the coefficients of $\mathcal{L}_{0}(\bar{z})$ and $\delta(\bar{z})$ in $d_{1\,Q/Q}^{(2)}$ at two loops.
-The required ingredients are readily available in the literature [44], [^4] The secondary mass effects in $\sqrt{S}$ have been calculated in ref.
+We stress that starting at $\mathcal{O}(\alpha_{s}^{2})$, the soft and collinear-soft functions do develop a highly nontrivial dependence on $mb_{T}$, which is thus inherited by the coefficients of $\mathcal{L}_{0}(\bar{z})$ and $\delta(\bar{z})$ in $d_{1\,Q/Q}^{(2)}$ at two loops.
+The required ingredients are readily available in the literature [74, 48], [^5] The secondary mass effects in $\sqrt{S}$ have been calculated in ref.
 [48].
@@
 7.1 Mass-dependent EEC jet functions
-In this section we use our results to produce explicit results for the the effect of heavy-quarks on the energy-energy correlator (EEC) observable [70] as measured in $e^{+}e^{-}$ collisions,
+In this section we use our results to produce explicit results for the effect of heavy quarks on the energy-energy correlator (EEC) observable [76] as measured in $e^{+}e^{-}$ collisions,
 $\mathrm{EEC}(\chi)\equiv\frac{1}{\sigma_{\mathrm{total}}}\,\frac{\mathrm{d}\sigma}{d\cos\chi}=\frac{1}{\sigma_{\mathrm{total}}}\sum_{a,b\,\in X}\int\!\mathrm{d}\sigma_{ee\to X}\frac{E_{a}E_{b}}{Q^{2}}\delta\bigl{[}\cos(\theta_{ab})-\cos(\chi)\bigr{]}\,,$ (7.1)
@@
 Specifically, in ref.
-[13] the characteristic denominator $1/X\equiv 1/(k_{T}^{2}+m^{2}\bar{z}^{2}/z^{2})$ appearing in the real-emission diagrams was directly expanded into one-dimensional plus distribution of $X$ and in addition was multiplied by one-dimensional distributions of $1-z$, which effectively amounts to performing a change of variables to $(z,X)$.
+[13] the characteristic denominator $1/X\equiv 1/(k_{T}^{2}+m^{2}\bar{z}^{2}/z^{2})$ appearing in the real-emission diagrams was directly expanded into one-dimensional plus distributions of $X$ and in addition was multiplied by one-dimensional distributions of $1-z$, which effectively amounts to performing a change of variables to $(z,X)$.
 This procedure is completely orthogonal to our approach, namely, we expand in terms of one-dimensional plus distributions along the boundaries $\bar{z}=0$ and $k_{T}=0$ and a two-dimensional plus distribution whose defining domain aligns with the main variables $(z,k_{T})$.
 The two approaches in particular differ (and have to differ) for the boundary contribution proportional to $\delta(1-z)$.
-(Indeed, the process in ref.
-[13] generates highly nontrivial twofold integrals $F(m^{2}/\Lambda^{2})$ and $G(m^{2}/\Lambda^{2})$ multiplying $\delta(1-z)$, where $\Lambda$ is the boundary condition of the $X$ plus distributions, with no closed form for $F$ and $G$ away from asymptotic limits for $m$.)
+Indeed, the process in ref.
+[13] generates highly nontrivial twofold integrals $F(m^{2}/\Lambda^{2})$ and $G(m^{2}/\Lambda^{2})$ multiplying $\delta(1-z)$, where $\Lambda$ is the boundary condition of the $X$ plus distributions, with no closed form for $F$ and $G$ away from asymptotic limits in $m$.
 In view of these substantial differences, it is highly nontrivial that the final results in $b_{T}$ space are indeed in full agreement, in particular regarding the latter’s $\delta(1-z)$ boundary condition.
@@
 The approach of ref.
-[13] checks the differential $z$ dependence of $d_{1\,Q/Q}^{(1)}$, but in intermediate steps an identical plus distribution expansion with respect to a very similar, complicated denominator $1/X^{\prime}=1/(k_{T}^{2}+\bar{z}^{2}m^{2})$, which is subject to the same subtleties as the full calculation.
+[13] checks the differential $z$ dependence of $d_{1\,Q/Q}^{(1)}$, but in intermediate steps an identical plus distribution expansion with respect to a very similar, complicated denominator $1/X^{\prime}=1/(k_{T}^{2}+\bar{z}^{2}m^{2})$ is used, which is subject to the same subtleties as the full calculation.
 In this sense, our calculation provides a strong complementary check of the large-mass behavior since it directly probes the total integral, i.e., a defining property of the distributional result.
@@
 Our stance here is that while – again – the assumption of ref.
-[13] is sufficient and entirely correct to describe specific regions of the multi-differential $(z_{H},k_{T})$ distribution, it is not the unique power-counting assumption on $1-z_{H}$, $k_{T}$, $m$ that would yield a leading-power factorization theorem involving bHQET modes, in particular not when different possible scalings of $\Lambda_{\mathrm{QCD}}$ are also considered.
+[13] is sufficient and entirely correct to describe specific regions of the multi-differential $(z_{H},k_{T})$ distribution, it is not the unique power-counting assumption on $1-z_{H}$, $k_{T}$, and $m$ that would yield a leading-power factorization theorem involving bHQET modes, in particular not when different possible scalings of $\Lambda_{\mathrm{QCD}}$ are also considered.
 As one example, we point to our analysis in section 6, where we continued to count $m\sim k_{T}$, but still took $1-z_{H}$ small, and likewise obtained a leading-power factorization theorem, in that case also involving a purely collinear bHQET shape function, but nontrivial $k_{T}$-dependent correlators of lightlike Wilson lines at the scale $\mu\sim k_{T}$.
-In light of this, we believe that capturing the full $(z_{H},k_{T})$ dependence will require dedicated future studies combining all of these results, while for the time being and within the scope of exploratory theory studies (or, necessarily less differential, initial experimental measurements) the restriction to wide bins or total integrals over $z_{H}$ is reasonable.
+In light of this, we believe that capturing the full $(z_{H},k_{T})$ dependence will require dedicated future studies combining all of these results, while for the time being and within the scope of exploratory theory studies (or, necessarily less differential, initial experimental measurements) the restriction to wide bins or low-moment integrals over $z_{H}$ is reasonable.
 9 Conclusions
@@
 Acknowledgments
-We thank Iain Stewart and Kyle Lee for helpful discussions.
+We thank Iain Stewart, Kyle Lee, and Maximilian Stahlhofen for helpful discussions.
 We thank the Erwin Schrödinger Institute for hospitality while parts of this work were performed.
@@
 Large-mass and joint-resummation limits.
-The combined collinear and soft mass matching coefficient $C_{m}$, see ref. [1] for our notation, is given by [43]
+The combined collinear and soft mass matching coefficient $C_{m}$ see also ref. [1] for our notation, is given by [47]
 $C_{m}(m,\mu,\zeta)=1+\frac{\alpha_{s}C_{F}}{4\pi}\biggl{(}4\ln^{2}\frac{\mu}{m}+2\ln\frac{\mu}{m}+4+\frac{\pi^{2}}{6}\biggr{)}+\mathcal{O}\bigl{(}\alpha_{s}^{2}\bigr{)}\,.$ (B.2)
-The $\zeta$ dependence arises from the rapidity renormalization of secondary soft heavy-quark contributions starting at $\mathcal{O}(\alpha_{s}^{2})$.
+At this order, it is also equal to the matching coefficient $H_{m,n}^{Q\to h_{v}}(m,\mu,\nu/\sqrt{\zeta})$ (defined in ref.
+[47], without the superscript) for matching massive collinear quark modes with label momentum $\sqrt{\zeta}$ onto bHQET on their own, whereas the additional contribution $\sqrt{H_{m,s}}(m,\mu,\nu)$ from matching secondary soft heavy quarks onto light soft modes becomes nontrivial at $\mathcal{O}(\alpha_{s}^{2})$.
+Two-loop expressions for both are given in ref.
+[47], where they were extracted from the NNLO heavy-quark form factor [79, 80].
+At that order, there is also a nonzero $\zeta$ dependence that arises from the rapidity renormalization of $H_{m,n}^{Q\to h_{v}}$ and $H_{m,s}$.
+The two-loop expression for secondary heavy-quark effects on the light-to-light collinear matching coefficient $H_{m,n}^{q\to q}$ is given in ref.
+[48] as $H_{c}$, where it was extracted from the results of refs.
+[81, 47].
 The perturbative HQET shape function is given by [5, 6]
@@
 The renormalized TMD soft function for both the $\eta$ regulator of ref. [43, 44] and the exponential regulator of ref. [71] reads
-$\sqrt{S}(b_{T},\mu,\nu)=1+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-L_{b}^{2}+4L_{b}\ln\frac{\mu}{\nu}-\frac{\pi^{2}}{6}\Bigr{]}\,.$ (B.4)
+$\sqrt{S}(b_{T},m,\mu,\nu)=1+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-L_{b}^{2}+4L_{b}\ln\frac{\mu}{\nu}-\frac{\pi^{2}}{6}\Bigr{]}\,.$ (B.4)
 The joint-resummation collinear-soft function is given by [72]
-$\mathcal{S}(\ell^{-},b_{T},\mu,\nu)=\delta(\ell^{-})+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-8L_{b}\,\mathcal{L}_{0}(\ell^{-},\nu)\Bigr{]}\,.$ (B.5)
+$\mathcal{S}(\ell^{-},b_{T},m,\mu,\nu)=\delta(\ell^{-})+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-8L_{b}\,\mathcal{L}_{0}(\ell^{-},\nu)\Bigr{]}\,.$ (B.5)
 Small-mass limit.
@@
 For $d_{1\,Q/Q}$, the relevant two-dimensional plus distribution simply turns into
-$\int_{0}^{1}\!\mathrm{d}z\,z^{N}\,\bigl{[}f(x,z)\bigr{]}_{+,+}=\Bigl{[}\int_{0}^{1}\!\mathrm{d}z\,\bigl{(}z^{N}-1\bigr{)}f(x,z)\Bigr{]}\,,$ (C.2)
+$\int_{0}^{1}\!\mathrm{d}z\,z^{N}\,\bigl{[}f(x,z)\bigr{]}_{+,+}=\Bigl{[}\int_{0}^{1}\!\mathrm{d}z\,\bigl{(}z^{N}-1\bigr{)}f(x,z)\Bigr{]}_{+}\,,$ (C.2)
 which like the moments of the one-dimensional plus distribution in $z$ can readily be evaluated for general $N$.
@@
 $\mathcal{J}_{1\,g/Q}^{[1]\,(1)}(k_{T},m,\mu,\zeta)$ $=\frac{\alpha_{s}C_{F}}{4\pi}\frac{1}{\pi m^{2}}\frac{2(1+4x)}{3(1+x)^{2}}\,.$ (C.3)
-Refer to captionRefer to caption
-Figure -1074: Left: Numerical Mellin moments $\mathcal{R}^{[N]}(y)$ of the finite remainder term in $d_{1\,Q/Q}$.
+Figure -1194: Left: Numerical Mellin moments $\mathcal{R}^{[N]}(y)$ of the finite remainder term in $d_{1\,Q/Q}$.
 Right: Numerical Mellin moments $d_{1\,Q/g}^{[N]\,(1)}$ of the gluon-to-massive-quark TMD FF.
@@
 We were not able to find a closed-form expression for the Mellin transform ~$\tilde{\mathcal{R}}^{[N]}(y)\equiv\int_{0}^{1}\!\mathrm{d}z\,z^{N}\,\tilde{\mathcal{R}}(z,y)$ of the finite remainder term, but implementing the $z$ integral of eq.
-(3.4) directly leads to stable numerical results (shown in the left panel of figure -1074 for reference) that could readily be interpolated.
+(3.4) directly leads to stable numerical results (shown in the left panel of figure -1194 for reference) that could readily be interpolated.
 For $\mathcal{J}_{g/Q}$, the Mellin transform for general $N$ is trivial since $z$ does not appear in the arguments of the Bessel $K$ functions,
@@
 (3.5).
-We show the latter in the right panel of figure -1074 for reference.
+We show the latter in the right panel of figure -1194 for reference.